### v1

# A Statistical Test for Comparing the Linkage and Admixture Model Based on Central Limit Theorems

Abstract
In the Admixture Model, the probability that an individual carries a certain allele at a specific marker depends on the allele frequencies in $K$ ancestral populations and the proportion of the individual’s genome originating from these populations. The markers are assumed to be independent. The Linkage Model is a Hidden Markov Model (HMM) that extends the Admixture Model by incorporating linkage between neighboring loci.
This study investigates the consistency and central limit behavior of maximum likelihood estimators (MLEs) for individual ancestry in the Linkage Model, complementing earlier results by (Pfaff et al., 2004; Pfaffelhuber and Rohde, 2022; Heinzel, 2025) for the Admixture Model. These theoretical results are used to prove theoretical properties of a statistical test that allows for model selection between the Admixture Model and the Linkage Model. Finally, we demonstrate the practical relevance of our results by applying the test to real-world data from The 1000 Genomes Project Consortium (2015).
Carola Sophia Heinzel
Department of Mathematical Stochastics,
Ernst-Zermelo Straße 1, Freiburg im Breisgau, 79140,Germany
carola.heinzel@stochastik.uni-freiburg.de
Keywords: Linkage Model, Admixture Model, Central Limit Results, Maximum Likelihood Estimator, Consistency, Statistical Test, Model Selection
1 Introduction
Both the Linkage Model (Falush et al., 2003) and the Admixture Model (e.g. Alexander et al. (2009)) are widely used to explain the genetic data of individuals. They are similar in that both assume that genetic data can be described by the ancestry proportions $q$ of an individual from $K$ ancestral populations and the allele frequencies in these populations. However, the Linkage Model additionally incorporates the distance between loci through a Hidden Markov Model (HMM). Furthermore, in the Linkage Model, there exists a parameter $r$, which can be interpreted as the number of generations since an admixture event (Falush et al., 2003). As a result, Falush et al. (2003) formulated the Linkage Model as a Hidden Markov Model (HMM). The Admixture Model can be seen as a special case of the Linkage Model with $r=\infty$, i.e. the data is assumed to be independent across markers.
A natural question that arises is which model fits a given dataset better. This leads to a nested model selection problem as described by Anderson and Burnham (2004). In our case, we consider the statistical hypothesis test:
$H_{0}:r=\infty\quad\text{vs.}\quad H_{1}:r\in[0,\infty).$ (1)
This is a classical nested test problem, for which asymptotic theory exists (Wilks, 1938). However, the test is only valid as an asymptotic level-$\alpha$ test if the MLEs for both ancestry and $r$ are asymptotically normally distributed under the Linkage Model. Establishing asymptotic normality requires results from the theory of HMMs.
Consistency of MLEs in HMMs with finite state and observation spaces has been addressed in foundational work by Baum and Petrie (1966) and Petrie (1969). More recent work has relaxed several of the original assumptions, e.g., (Leroux, 1992; Douc et al., 2004, 2011; Genon-Catalot and Laredo, 2006; Le Gland and Mevel, 2000a, b), but these results assume time-homogeneous Markov chains. Central limit theorems for MLEs in HMMs have also been developed (Bickel et al., 1998; Douc et al., 2004; Jensen and Petersen, 1999; Brouste and Kleptsyna, 2010). Overviews of the statistical theory of HMMs can be found in Ephraim and Merhav (2002); Cappé et al. (2005). However, the specific case of a time-inhomogeneous Markov chain that maintains the same stationary distribution across all time points—as occurs in the Linkage Model when $q$ is the initial distribution—has not yet been studied.
Several statistical tests and heuristics have been proposed for model selection in genetics. For example, to decide whether a pair of loci is suitable for STRUCTURE (Pritchard et al., 2000), Kaeuffer et al. (2007) suggested using $r_{LD}$, a measure of linkage disequilibrium introduced by Hill and Robertson (1968). They ran STRUCTURE with the Linkage Model and used logistic regression to assess the impact of $r_{LD}$ on the detection of population structure. Other studies have examined the sensitivity of cluster identification to genetic and ecological constraints (Rosenberg et al., 2001; Evanno et al., 2005).
To achieve our goal of establishing the theoretical properties of the test in (1), we proceed as follows: First, we precisely define both the Admixture Model and the Linkage Model. Next, we prove the asymptotic normality of the MLEs in the Linkage Model, both when the number of chromosomes and when the number of markers tends to infinity. Based on these results, we construct a statistical test for (1) and prove that it is an asymptotic level-$\alpha$ test. We also provide a method to quantify the uncertainty of the MLEs in the Linkage Model. Finally, we evaluate the statistical test through simulations and apply it to real data from The 1000 Genomes Project Consortium (2015).
2 Models
We first define the Linkage Model for one individual and bi-allelic markers.
Definition 2.1 (Linkage Model for Haploid Individuals).
The number of alleles at chromosome $c\in\{1,...,C\}$ at marker $m\in\{1,...,M_{c}\}$ is called $X_{c,m}\in\{0,1\}.$ We denote the genetic distance in centi Morgan (cM) between the loci m1m-1 and $m$ on chromosome $c$ by $d_{c,m}$. Let $\mathbb{S}^{K}$ be the (K1)(K-1)-dimensional simplex and let $q:=(q_{1},\ldots,q_{K})\in\mathbb{S}^{K}$ be the ancestries of the individual from population 1,,K.1,...,K. The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, chromosome $c$ is called $p_{c,k,m}.$ The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ at chromosome $c$. Based on this, we define the Markov process
$\mathbb{P}_{q,r}(Z_{c,1}=k)$ $=q_{k},$
$\mathbb{P}_{q,r}(Z_{c,m}=\tilde{k}|Z_{c,m-1}=k)$ $=\begin{cases}e^{-d_{c,m}r}+\left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ if }k=\tilde{k}\\ \left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ else},\end{cases}$ (2)
$\mathbb{P}_{q,r}(Z_{c+1,1}=k|Z_{c,M_{c}})$ $=q_{k}.$
The emission probability is defined by
$\mathbb{P}_{q,r}(X_{c,m}=x|Z_{c,m})=\mathbb{P}_{q,r}\left(\textup{Ber}\left(q_{Z_{c,m}}p_{c,Z_{c,m},m}\right)=x\right).$
We define the log-likelihood, for $M_{total}:=\sum_{c=1}^{M_{C}},$
$\ell((x_{1,1},...,x_{C,M}),(q,r)):=\frac{\log\left(\mathbb{P}_{q,r}((X_{c,m}=x_{c,m})_{c=1,...,C,m=1,...,M_{c}}))\right)}{M_{total}}.$
We write $(q^{0},r^{0})$ for the true parameters $(q,r)$ and $\mathbb{E},\mathbb{P}$ for the expected value and the probability respectively, with respect to the true parameters $q^{0},r^{0}.$
We always assume that $d_{c,m}\geq 0$ for all $m\in\{1,\ldots,M\}$ and that the allele frequencies and $d_{c,m}$ are known.
Remark 2.2 (Diploid Case).
Let $Z^{j}:=(Z^{j}_{c,1},...,Z^{j}_{c,M}),j=1,2,$ be two independent Markov chains with transition matrix (2). Let $X_{c,m}^{dip}$ be the number of alleles at chromosome $c,$ marker $m$ and let $x\in\{0,1,2\}.$ The emission probabilities for the diploid case are defined by
$\mathbb{P}_{q,r}\left(X^{dip}_{c,m}=x|Z^{1}_{c,m},Z^{2}_{c,m}\right)$
$\hskip 28.45274pt=\begin{cases}q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}\cdot q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m},&x=2\\ q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}(1-q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m})+q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m}(1-q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}),&x=1\\ (1-q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m})(1-q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m}),&x=0.\end{cases}$
For the diploid case, the information about the maternal and the paternal copies (Choi et al., 2018), i.e. phased data, is important. There are two kinds of approaches to receive phased data: laboratory-based methods (Zheng et al., 2016; Amini et al., 2014; Duitama et al., 2012) and computational methods (Choi et al., 2018). Examples for the latter type are (Snyder et al., 2015; Loh et al., 2016; Delaneau et al., 2012). Falush et al. (2003) also proposed a method to deal with linkage without phasing. However, in this study, we assume that we have phased data as e.g. in The 1000 Genomes Project Consortium (2015).
While phased data is important for the Linkage Model, the Admixture Model can deal with unphased data without any problems.
Definition 2.3 (Admixture Model).
The Admixture Model is a special case of the Linkage Model with $r=\infty.$
Extending the models to a general number of individuals and markers with arbitrary number of alleles is straightforward.
We do not only need to know about the Linkage Model, but we also need some general notation is mentioned in Notation 2.4.
Notation 2.4 (General Notation).
For a positive definite matrix $A$, we write $A\succ 0.$ Additionally, we write $X_{i:i+s}$ for $X_{i},...,X_{i+s}.$
3 Theoretical Main Results
In this chapter, we prove the main results, i.e. consistency and central limit results, if the number of markers, $M_{total},$ tends to infinity. Therefore, we need assumption 3.1.
Assumption 3.1.
We assume that
  • (A1)
    the parameter space $\Theta:=\{[\kappa_{q},\kappa_{q}^{\prime}]^{K-1}\times[r_{lb},\infty]\}\subset\mathbb{R}^{K+1}$ is a compact set.
  • (A2)
    it holds $1>\kappa_{p}^{\prime}\geq p_{\cdot,m}\geq\kappa_{p}>0$ for all $m\in\{1,...,M\}.$
  • (A3)
    $p_{k,m}\neq p_{\ell,m}$ for infinitely many markers m and for all $k,\ell\in\{1,..,K\},k\neq\ell.$
  • (A4)
    Let $\lambda_{c,m}$ be the second largest of the transition matrix $(\mathbb{P}(Z_{c,m}=k|Z_{c,m-1}=k^{\prime}))_{k=1,...,K,k^{\prime}=1,...,K}$. It holds
    $\prod_{c=1}^{C}\prod_{m=1}^{M_{c}}\lambda_{c,m}\xrightarrow[]{M_{total}\to\infty}0$
  • (A5)
    There exists a constant $\kappa_{d}>0$, such that $d_{c,m}\geq\kappa_{d}$ for all $m\in\{1,...,M_{c}\},c=1,...,C.$
Figure 1 gives an overview of the results in this paper.
ConsistencyCLTAsymptotic Distribution of the Test Statistic
Figure 1: Overview of results in this section and how they are connected to each other.
We also state the theory under the following conditions concerning the data.
Remark 3.2 (Assumptions).
We assume that the data is haploid and that we only have bi-allelic markers. Additionally, we only consider one individual. However, extending the theory to more general cases is straightforward.
We start with the consistency of the MLE (Theorem 1).
Theorem 1 (Consistency of the MLE).
Let Assumption 3.1 hold. Then, for the MLE
$\left(\hat{Q}^{C,M},\hat{R}^{C,M}\right):=\textup{argmax}\{(q,r)\mapsto\ell((x_{1,1},...,x_{C,M}),(q,r))\}$
it holds
$\mathbb{P}\left(\lim_{M_{total}\to\infty}\bigg{|}\left(\hat{Q}^{C,M},\hat{R}^{C,M}\right)-\left(q^{0},r^{0}\right)\bigg{|}\geq\epsilon\right)=0$
for any $\epsilon>0.$
Based on Theorem 1, we can infer a CLT for the MLE, if the true parameters $\left(q^{0},r^{0}\right)$ are in the interior of the parameter space.
Theorem 2 (Central Limit Theorem for the MLE).
We define the Fischer information
$J_{q^{0},r^{0}}:=\lim_{M_{total}\to\infty}\mathbb{E}\left(\frac{\partial^{2}}{\partial(q,r)^{2}}\ell\left((x_{1,1},...,x_{C,M}),(q,r))\big{|}_{(q,r)=\left(q^{0},r^{0}\right)}\right)\right)$
and assume $J_{q^{0},r^{0}}\succ 0$. Let the true parameter be in the interior of the parameter space. Then, it holds
$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow{M_{total}\to\infty}\mathcal{N}\left(0,J_{q^{0},r^{0}}^{-1}\right).$
Remark 3.3.
Douc (2005) has proven the invertibility of the Fischer Information under certain constraints for stationary HMMs, i.e. in our case, if the allele frequencies and the distance between the markers are identical for every marker. This also means that the markers are all at different chromosomes or at the same chromosome.
According to Wilks (1938), since the MLE in the Linkage Model is asymptotically normally distributed, Theorem 3 holds.
Definition 3.4 (Statistical Test).
We define the test statistic for the test (1) by
$\Lambda:=-2\ln\left(\frac{\textup{argmax}\{(q,r)\mapsto\ell((X_{1,1},...,X_{C,M}),(q,\infty))\}}{\textup{argmax}\{(q,r)\mapsto\ell((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
Furthermore, let $\chi^{2}_{1-\alpha}$ be the $1-\alpha$-quantile of the $\chi^{2}(1)$-distribution. We reject $H_{0}$, if $\Lambda>\chi^{2}_{1-\alpha}.$
Theorem 3 (Asymptotic distribution of the test statistic).
It holds
$\Lambda\xRightarrow{M_{total}\to\infty}\chi^{2}(1)$
under the null hypothesis of (1).
4 Application to Data
In this chapter, we first evaluate the performance of the test from Definition 3.4 by using simulated data. Afterwards, we apply the test to data from The 1000 Genomes Project Consortium (2015) and compare the uncertainty of the MLEs in the Linkage Model to the ones in the Admixture Model.
4.1 Simulated Data
To evaluate the statistical test by using the type 1 and type 2 error, we simulate the data according to the Linkage Model and the Admixture Model with different values for $M,r$ and $(d_{c,m})_{c=1,...,C,m=1,...,M_{c}-1}.$ More precisely, we simulated 100 markers on one single chromosome with $d_{1,m}=d_{1,m+1}\forall m,d:=d_{1,m}\in\{0.1,0.5,1,2,5,10\}$ and $r\in\{1,10,100\}.$ We choose $0.05$ as the significance level. We calculated the type 1 and the type 2 error by repeating the experiment 100 times per possible combination of $r$ and $d$.
The results are shown in Figure 2. Especially the type 1 error is small and lower than 0.05. Furthermore, for increasing $r$ and $d$ the type 1 and the type 2 error tend to increase. This is not surprising since a large $r$ corresponds to a high similarity between the two models. To summarize, the simulation results show that the test has both, a high power and a small type 1 error.
Figure 2: Evaluation of the statistical test for simulated data.
4.2 Real Data
We also applied our theoretical results to the data from The 1000 Genomes Project Consortium (2015) with the AIM set by Kidd et al. (2014) that consists of $M_{total}=55$ markers on 20 different autosomes ($C=20$). We always used all 2504 individual except the one that we are testing to calculate the allele frequencies as relative frequency. The values for $d$ for the AIM set by Kidd et al. (2014) in cM are shown in Figure 3.
Figure 3: Genetic Distances of the markers in the AIM set by Kidd et al. (2014). We rounded them to integer values.
The distance between the loci at a single chromosome was between below 1 cM and approximately 129 cM. We used individuals from Africa (AFR), Europe (EUR), South-East-Asian (SAS), East-Asian (EAS) and Admixed Americans (AMR), i.e. $K=5$.
The results of the statistical test are shown in Figure 4. Altogether, for 312 out of 2504 ($\approx 12.5\%$ of the) individuals, the null hypothesis that the Admixture Model fits better to the data, cannot be rejected. Thereby, we see strong differences between the populations: For people from Africa and East Asian, the fraction of individuals for which the null hypothesis is rejected is much smaller than for the other populations. This emphasizes that it is not sufficient to name a general rule to decide for either the Linkage Model or the Admixture Model that only depends on the marker set. Instead, a method – like the statistical test (1) – is required, that decides for every individual or population separately.
Figure 4: Results of the statistical test.
We also compare the covariance for both, the MLE in the Linkage Model and the MLE in the Admixture Model to each other for individual HG00096 from Great Britain. Therefore, we again used the AIM Set by Kidd et al. (2014). The results are presented in figures 6 and 6. There, we see that the variance of the estimators in the Linkage Model are higher than in the Admixture Model. This aligns with our expectations since the Linkage Model contains more parameters. Although both models clearly identify EUR as the primary ancestral population, the estimated values differ significantly.
Figure 5: Covariance Matrix for the MLE in the Admixture Model. We considered individual HG00096. The MLE for $q$ was $(1,0,0,0,0).$
Figure 6: Covariance Matrix of the MLE in the Linkage Model. We considered individual HG00096. The MLE for $q$ was $(0.8,0.04,0.04,0.08,0.04)$ and for $r$ approximately $1.8.$
Remark 4.1 (Test for a whole population).
We also provide the code for testing whether the Linkage Model or the Admixture Model fits better to the data of $N$ individuals on GitHub. Therefore, we could either assume that $r$ depends on the individual, i.e. we would have $NK$ parameters that we have to estimate or that we have only one global parameter $r,$ i.e. we estimate N(K1)+1N(K-1)+1 parameters. We decided for the latter case. However, the allele frequencies are supposed to be known in this case. Additionally, the diploid case is also implemented.
5 Proofs of the Main Results
We first prove that the MLE is asymptotically unique. Based on this, we prove consistency and central limit results for the MLE in the Linkage Model. For all proofs, we assume $C=1$ to simplify the notation.
To prove the consistency and the CLT, we represent the log-likelihood as a sum, i.e. it holds
$\ell((X_{1},...,X_{M}),(q,r))$
$\quad=\frac{1}{M}\sum_{m=1}^{M}\underbrace{\log\left(\int\mathbb{P}_{q,r}(X_{m}|Z_{m}=z)\mathbb{P}_{q,r}(Z_{m}=dz|X_{1},...,X_{m-1})\right)}_{=:D_{m}^{q,r}}$
according to van Handel (2008), Proposition 6.4, for the homogeneous case. This can easily extended to the inhomogeneous case.
5.1 Unique MLE
In this section, we use that the limit
$\ell((X_{1},...,X_{M}),(q,r))\xrightarrow[]{M\to\infty}\ell(q,r)$
exists almost surely. We will prove this in section 5.2.
Proving the uniqueness of a MLE for HMM has been already considered extensively for a finite state space, see e.g. (Finesso, 1990; Gilbert, 1959; Blackwell and Koopmans, 1957; Petrie, 1969). In this section, we aim to prove the asymptotic uniqueness of the MLE in the Linkage Model, i.e. we aim to prove Theorem 4.
Theorem 4.
Let Assumption 3.1 hold. Then, it holds
$\ell(q^{0},r^{0})>\ell(q,r)\Leftrightarrow q,r\neq q^{0},r^{0}.$ (3)
Therefore, we first prove that the parameters are identifiable based on a finite number of observations. Based on this, we prove that if $\ell(q^{0},r^{0})\neq\ell(q,r)$, then the measures $\mathbb{P}_{q,r}|_{(X_{m})_{m=1,2,...}}$ and $\mathbb{P}_{q^{0},r^{0}}|_{(X_{m})_{m=1,2,...}}$ are mutually different. Based on this, we can then conclude the claim.
We write $(q,r)\sim(q^{\prime},r^{\prime})$, if there exists an invariant distribution $\pi,\pi^{\prime}$ such that $\mathbb{P}_{(q,r)}^{X}=\mathbb{P}_{(q^{\prime},r^{\prime})}^{X}$ holds. Therefore, we only consider the case $K=2$ as the proof involves naming conditions that yield to the full rank of the emission matrix.
Lemma 5.1 (Identifiability of the Parameters).
Let Assumption 3.1 hold and let $K=2.$ If either the allele frequencies $p_{\cdot,i},...,p_{\cdot,i+s}$ are linearly independent and $s\geq 4$ or $s=3$ and $q\neq p_{1,2}/(p_{1,2}+p_{2,2})$ holds and in addition $p_{1,m}\neq p_{2,m}\forall m\in\{i,...,i+s\}$, the parameters $(q,r)$ are identifiable based on the random variables $X_{i},...,X_{i+s}.$
Proof.
To save some notation, we write $A_{i}(z_{i},z_{i+1})=\mathbb{P}(Z_{i+1}=z_{i+1}|Z_{i}=z_{i}),B_{j}(x_{j},z_{j})=\mathbb{P}(X_{j}=x_{j}|Z_{j}=z_{j}).$ In order to apply Kruskal’s Theorem, we first need a decomposition of the marginal density of $X_{1},X_{2},X_{3}.$ Let $M_{j}$ be a stochastic matrix, where the $i^{th}$ row is denoted by $m^{j}_{i}.$ We define
$[M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}=\sum_{i=1}^{2}m^{1}_{i}(x_{1})m^{2}_{i}(x_{2})m^{3}_{i}(x_{3})$
and $M_{1},M_{2},M_{3}]$ is a three-dimensional array for which the element $(x_{1},x_{2},x_{3})$ equals $M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}.$ It holds
$\mathbb{P}(X_{1},X_{2},X_{3})$ $=\sum_{z_{1},z_{2},z_{3}}\pi_{z_{1}}A_{1}(z_{1},z_{2})A_{2}(z_{2},z_{3})B_{2}(X_{2},z_{2})B_{3}(X_{3},z_{3})B_{1}(X_{1},z_{1})$
$=[\underbrace{\pi B_{1}A_{1}}_{=:M_{1}},B_{2},\underbrace{A_{2}B_{3}}_{=:M_{2}}].$
The proof is divided into two steps:
  • (i)
    prove that the entries of $M_{1},M_{2},B_{2}$ are identifiable from the marginal density of $X_{1},X_{2},X_{3}$. Therefore, we use theory by Allman et al. (2009); Kruskal (1977).
  • (ii)
    conclude that the parameters $q,r$ are identifiable.
We start with (i). Therefore, we have to prove $rk(M_{1})=rk(B_{2})=rk(M_{2})=2.$ It holds
$\det(B_{2})=\bigg{|}\begin{pmatrix}q_{1}p_{1,2}&(1-q_{1})p_{2,2}\\ 1-q_{1}p_{1,2}&1-(1-q_{1})p_{2,2}\end{pmatrix}\bigg{|}=q_{1}p_{1,2}-(1-q_{1})p_{2,2}.$
Hence, $B_{2}$ has the full rang if and only if $q_{1}\neq\frac{p_{1,2}}{p_{1,2}+p_{2,2}}.$ This is the reason, why we need the condition about the linear independence of the allele frequencies (since we can apply Kruskal’s theorem to two different time points within the interval). We have
$M_{1}=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$
with entries
$a$ $=(1-q_{1}p_{1,1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+\left(1-(1-q_{1})p_{2,1}\right)(1-e^{-rd_{1}})(1-q_{1})$
$=(1-q_{1}p_{1,1})(q_{1}+e^{-rd_{1}}(1-q_{1}))+(1-p_{2,1}+q_{1}p_{2,1})(1-q_{1})(1-e^{-rd_{1}})$
$b$ $=q_{1}p_{1,1}\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+p_{2,1}(1-q_{1})^{2}(1-e^{-rd_{1}}),$
$c$ $=(1-q_{1}p_{1,1})(1-e^{-rd_{1}})(1-q_{1})+\left(1-(1-q_{1})p_{2,1}\right)\left((1-e^{-rd_{1}})(1-q_{1})+e^{-rd_{1}}\right),$
$d$ $=q_{1}p_{1,1}(1-e^{-rd_{1}})(1-q_{1})+p_{2,1}(1-q_{1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right).$
It holds
$\det(M_{1})=e^{-rd_{1}}(1-e^{-rd_{1}})q_{1}(1-q_{1})(p_{1,1}+p_{2,1}-2p_{1,1}p_{2,1})=0$
$\Leftrightarrow(r=\infty)\lor(r=0)\lor(q_{1}\in\{0,1\})\lor(p_{1,1}=p_{2,1}=0)\lor(p_{2,1}=p_{1,1}=1).$
Here, we excluded every case except the case $r=\infty.$ This case leads to the Admixture Model which has already been discussed by Heinzel (2025). We additionally calculate
$\mathbb{P}(X_{k-2}=x\mid Z_{k-1}=z)$ $=\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x,Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
$\propto\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
$\propto\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-1}=z^{\prime}\mid Z_{k-2}=z)$
$\propto\underbrace{\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=0)}_{=(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}}\cdot\mathbb{P}(Z_{k-1}=0\mid Z_{k-2}=z)$
$\hskip 28.45274pt+\underbrace{\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=1)}_{=(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}}\cdot\underbrace{\mathbb{P}(Z_{k-1}=1\mid Z_{k-2}=z)}_{=1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})}$
$\propto(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}\left(1_{z=0}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right)$
$\hskip 28.45274pt+(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}\left(1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right).$
This leads to
$M_{2}=\begin{pmatrix}a_{2}&b_{2}\\ c_{2}&d_{2}\end{pmatrix}$
with entries
$a_{2}$ $=(1-q_{1}p_{1,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right)+(1-(1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}})$
$b_{2}$ $=(q_{1}p_{1,3})\left(e^{-rd_{2}}+q_{1}(1-e^{-rd_{2}})\right)+((1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}}),$
$c_{2}$ $=(1-q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})+(1-(1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right),$
$d_{2}$ $=(q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})+((1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right).$
Hence, it holds $rk(M_{2})=1\Leftrightarrow p_{2,3}=p_{1,3},q_{1}=0.5.$ Now, we come to step (ii), i.e. we have to ensure that the parameters are uniquely determined based on the entries. For the emission probabilities, this is a direct consequence of the assumption $p_{k,m}\neq p_{\ell,m}.$ Then, since $d_{m}>0,$ it is also immediate to infer the value $r$ from the other parameters.
Remark 5.2 (Identifiability in the Admixture Model for $K=2$).
Heinzel (2025) did not consider the identifiability for finitely many markers in the Admixture Model. For $K=2$, it holds
$\mathbb{P}(X_{m}=x)$ $=\langle q,p_{\cdot,m}\rangle^{x}(1-\langle q,p_{\cdot,m}\rangle)^{1-x}$
$=(q_{1}p_{1,m}+(1-q_{1})p_{2,m})^{x}(1-q_{1}p_{1,m}-(1-q_{1})p_{2,m})^{1-x}$
$=\begin{cases}p_{2,m}+q_{1}(p_{1,m}-p_{2,m})=:c_{m},x=1\\ 1-p_{2,m}+q_{1}(-p_{1,m}+p_{2,m}),x=0\end{cases}.$
This leads to
$q_{1}=\frac{c_{m}-p_{2,m}}{p_{1,m}-p_{2,m}},$
if $p_{1,m}=p_{2,m}.$ Hence, $q_{1}$ is uniquely determined if and only if $p_{1,m}=p_{2,m}.$
Remark 5.3.
In the latter, we will – without a proof – just assume that the parameters are identifiable from the marginal density of the observations. The proof is basically the same. However, we of course have to increase the number of observations form $3$ for $K=2$ to 2(K1)+2.2(K-1)+2. Additionally, we then have to verify again that the matrices of the decomposition of the marginal density have the full rank under certain conditions.
Now, we aim to infer (3) from the identifiability of the parameters. Therefore, we first prove a law of large numbers. Then, based on this, we conclude (3). This is done by using the following particularity of the hidden chain (Lemma 5.4) and that $(Z_{m})_{m=1,2,...}$ is uniformly ergodic (Lemma 5.6).
Lemma 5.4 (Invariant Measure).
For each $m\in\mathbb{N},$ the probability measure $q$ is the unique invariant measure for the transition matrix of $Z_{m}$.
Definition 5.5 (Uniformly Ergodic).
Let $|\cdot|_{TV}$ be the total variation norm. A Markov chain with transition matrices $T_{i}^{\theta}$ is called uniformly ergodic, if it holds
$|T^{\theta}_{n}...T^{\theta}_{1}-\pi|_{TV}\leq c_{\theta}\rho_{\theta}^{n}$
for $\rho_{\theta}<1,c_{\theta}\in\mathbb{R}.$
Lemma 5.6 (Uniformly Ergodicity for the Linkage Model).
Under Assumption 3.1, $(Z_{m})_{m=1,2,...}$ is uniformly ergodic.
Proof.
This claim is derived from Theorem 3.3 in Saloff-Coste and Zúniga (2007). ∎
Now, we are ready to name an LLN for the Linkage Model.
Lemma 5.7.
Let Assumption 3.1 hold and let $f_{i}$ be uniformly bounded by $f_{\infty}$. Then, it holds
$\mathbb{P}^{\nu}_{\theta}\left(\frac{1}{n}\left|\sum_{i=1,...,n}f_{i}(X_{i:i+s})-\pi(f_{i})\right|\geq\epsilon\right)\leq C\exp\left(-\frac{n}{8f_{\infty}}\right)$
for a constant $C>0$ and for any probability measure $\nu.$
Proof.
Without loss of generality, we assume $\mathbb{E}_{\pi}(f_{i}(X_{i:i+s}))=0$. From Lemma 5.6, we know that the Markov chain $(Z_{m})_{m=1,2,...}$ is uniformly ergodic. The proof is similar to Theorem 14 in Douc (2005): We first write
$\xi_{i,j}:=\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i+j},X_{0:i+j})+\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i+j-1},X_{0:i+j-1}).$
Then, since
$\sum_{j=0}^{s}\sum_{i=1}^{n}\xi_{i,j}$ $=\sum_{i=1}^{n}\sum_{j=0}^{s}\xi_{i,j}=\sum_{i=1}^{n}\xi_{i,s}-\xi_{i,0},$
it holds
$\sum_{i=1,...,n}f_{i}(X_{i:i+s})=\sum_{j=0}^{s}\sum_{i=1}^{n}\xi_{i,j}+\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i-1},X_{0:i-1}).$
Now, we can apply Azuma-Hoeffding inequality, which leads to
$\mathbb{P}^{\nu}\left(\bigg{|}\sum_{i=1}^{n}\xi_{i,j}\bigg{|}\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{8nf_{\infty}}\right).$
Now, we note that $\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|X_{0:i-1},Z_{0:i-1})=F(Z_{i-1})$ for a certain function $F.$ We now prove that
$\mathbb{P}\left(\bigg{|}\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|X_{0:i-1},Z_{0:i-1})\bigg{|}\geq t\right)2\leq\exp\left(-\frac{t^{2}}{8nf_{\infty}}\right).$
This is Theorem 17 in Douc et al. (2011).
To prove this claim, we apply Etemadi’s inequality and Bernstein’s inequality to the $\xi_{k}$ as defined in Lemma 18 in Douc et al. (2011). Neither of them requires that the $\xi_{k}$ are identically distributed. Additionally, they used their Proposition 19 to prove the claim. Note, however, that in our case we can highly simplify the proof as we have a discrete Markov chain with bounded transition probabilities. Especially, we do not even need their $\check{X}_{n}$ as in our case, their $d_{n}$ equals 1 for all $n\in\mathbb{N}.$
Lemma 5.8.
Assumption 3.1 implies that there exists a uniformly bounded function $h_{i}$ and an integer $s\in\mathbb{N}$ such that
$(i)\quad\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{M\to\infty}1\quad\mathbb{P}_{q^{0},r^{0}}-a.s.,$
$(ii)\quad\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{M\to\infty}0\quad\mathbb{P}_{q,r}-a.s.$
for every $(q,r)\neq(q^{0},r^{0}).$
Proof.
The proof is based on Lemma 6 in Douc et al. (2011). However, in our case $X$ is not a stationary sequence. We find an integer $s$ and bounded functions $h_{i},i=1,2,$..., such that $\mathbb{E}_{\pi}(h_{i}(X_{i},...,X_{i+s}))=1,\mathbb{E}_{\pi^{0}}(h_{i}(X_{i},...,X_{i+s}))=0$ according to the theorem from Hahn-Banach for $\pi\neq\pi^{0}$. This is a consequence of Lemma 5.1. More precisely, we choose
$h_{i}(X_{i},...,X_{i+s})=1-\frac{f_{i}(X_{i},...,X_{i+s})-1}{\mathbb{E}_{\pi^{0}}(f_{i}(X_{i},...,X_{i+s}))-1}$
with
$f_{i}(X_{i},...,X_{i+s}):=\frac{\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})}{\mathbb{P}_{\pi}(X_{i},...,X_{i+s})}$
for $\pi\neq\pi^{0}.$ The case $\mathbb{E}_{\pi^{0}}(f_{i})=1$ cannot occur, which follows from Cauchy-Schwarz. However, the densities are uniformly bounded according to Assumption 3.1, i.e. we can also choose $h_{i}$ uniformly bounded.
Hence, we use Dolgopyat and Sarig (2023), Theorem 3.12, to prove (i).
For (ii), We define
$A_{M}:=\left\{X_{0:M}:\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})\geq\frac{1}{2}\right\}.$
It holds
$\mathbb{P}_{q,r}^{\nu}(X_{0:M}\in A_{M})$
$\quad=\sum_{k=1}^{K}\mathbb{P}_{q,r}^{\nu}(X_{0:M}\in A_{M}|Z_{0}=k)\mathbb{P}_{q,r}^{\nu}(Z_{0}=k)$
$\quad\leq\max_{k=1,...,K}\sup_{\textup{supp }\mu}\mathbb{P}^{\mu}_{q,r}\left(\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})\geq\frac{1}{2}\right)$
$\quad\leq c_{1}\exp(-M/c_{2})$
for some constant $c_{1},c_{2}>0.$ Here, we use Lemma 5.7 for the last inequality. ∎
Finally, we use Lemma 5.8 to prove Theorem 4.
Proof of Theorem 4.
The direction $\Rightarrow"$ is trivial. The other direction is similar to van Handel (2008), Theorem 7.13. First, we prove
$\ell(q,r)=\ell(q^{0},r^{0})\Rightarrow\mathbb{P}_{q^{0},r^{0}}|_{(X_{m})_{m=1,2,...}}<<\mathbb{P}_{q,r}|_{(X_{m})_{m=1,2,...}}$.
Since the number of observable states is finite, we immediately obtain this.
Now, we prove that $\mathbb{P}_{q^{0},r^{0}}\neq\mathbb{P}_{q,r}$ implies that $\mathbb{P}_{q^{0},r^{0}},\mathbb{P}_{q,r}$ are mutually singular. This is Lemma 5.8. Hence, we can then infer the claim analogously to van Handel (2008) in the proof of Theorem 7.13. ∎
5.2 Consistency
We prove Theorem 1, i.e. the consistency of the MLE. Therefore, we follow the ideas by van Handel (2008), which are stated in the homogeneous case. First, we note that both the transition probabilities and the emission probabilities are Lipschitz continuous in the parameters given Assumption 3.1. This is a direct consequence of the mean value theorem applied to $r\mapsto|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|.$
We already know that the true value is the unique maximum point of the likelihood according to Theorem 4. Therefore, we first prove
$\sup_{(q,r)\in\Theta}|\ell((X_{1},...,X_{M}),(q,r))-\ell(q,r)|\xrightarrow[]{M\to\infty}0(*).$
Based on this, we almost immediately get the claim of Theorem 1.
Proving (*) is Lemma 5.11 whose proof is divided into three steps:
  1. 1)
    Prove $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$ Since both, the transition probabilities and the emission probabilities are uniformly bounded below, we can infer (according to the majorant criterion)
    $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$
  2. 2)
    Prove $\ell((X_{1},...,X_{M}),(q,r))$ convergences for $M\to\infty$ a.s. We call the limit $\ell(q,r).$ This is Lemma 5.9
  3. 3)
    Prove the claim ()(*). This is Lemma 5.10.
Lemma 5.9 is Lemma 7.8 in van Handel (2008), which easily can be adapted to the inhomogeneous case.
Lemma 5.9.
Let Assumption 3.1 hold and let
$D_{k,\ell}^{q,r}:=\log\left(\int\mathbb{P}_{q,r}(X_{k}|Z_{k}=z)\mathbb{P}_{q,r}(Z_{k}=z|X_{k-1},...,X_{k-\ell})dz\right)$
There exist constants $\epsilon\in(0,1),\alpha$ so that
$\sup_{k\in\mathbb{N}}|D_{k,\ell}^{q,r}-D_{k}^{q,r}|\leq\alpha(1-\epsilon^{2})^{\ell}$
for all $\ell,k\in\mathbb{N}.$
Lemma 5.10 is Lemma 7.9 in van Handel (2008).
Lemma 5.10.
Let Assumption 3.1 hold. There exists a constant $\alpha$ so that
$\sup_{k\in\mathbb{N}}|D_{k}^{q,r}-D_{k}^{q^{\prime},r^{\prime}}|\leq\alpha|(q,r)-(q^{\prime},r^{\prime})|_{2}$
for all $(q,r),(q^{\prime},r^{\prime})\in\Theta.$
Lemma 5.11.
Let Assumption 3.1 hold. Then, it holds
$\sup_{q,r\in\Theta}|\ell_{M}(q,r)-\ell(q,r)|\xrightarrow[]{M\to\infty}0.$
Proof.
For simplicity, we write $\theta=(q,r)$ and $\theta^{0}=\left(q^{0},r^{0}\right).$ We start chronologically:
  1. 2)
    We apply Lemma 7.7 in van Handel (2008) to $D_{k}^{\theta}-\mathbb{E}_{\theta^{0}}(D^{\theta}_{k})$, for which we have to check whether there exists $\rho\in(0,1),C\in\mathbb{R}$ so that
    $\mathbb{E}_{\theta^{0}}\left(D_{k}^{\theta}-\mathbb{E}_{\theta^{0}}(D^{\theta}_{k})|X_{1},...,X_{\ell}\right)\leq C\rho^{k-\ell}$
    holds. Therefore, we use Lemma 5.9.
  2. 3)
    The details can be found in van Handel (2008). Let us mention that we first prove that $\ell(q,r)$ is also Lipschitz continuous. Then, the claim is a direct consequence of Lemma 5.10 and the compactness of $\Theta$.
Proof of Theorem 1.
We first show that if the likelihood has a unique maximum in $(q^{0},r^{0})$, the MLE tends to this. Therefore, we calculate $0\leq\ell(q,r)-\ell(\hat{Q}^{M},\hat{R}^{M})\xrightarrow[]{M\to\infty}0$ according to Lemma 5.11. Additionally, we know that the MLE is unique, which leads to the claim. The details are described by van Handel (2008). ∎
5.3 Central Limit Results
To prove Theorem 2, we first prove that the first derivative of $\ell((X_{1},...,X_{M}),(q^{0},r^{0}))$ is asymptotically normally distributed.
Proposition 5.12.
Let $J(q^{0},r^{0})\succ 0$ and let Assumption 3.1 hold. Then, it holds
$\lim_{M\to\infty}\frac{\nabla\ell((X_{1},...,X_{M}),(q^{0},r^{0}))}{\sqrt{M}}\xRightarrow{M\to\infty}\mathcal{N}\left(0,J^{-1}\left(q^{0},r^{0}\right)\right).$
Proof.
Let $H^{(q^{0},r^{0})}_{k}:=\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)$. First, we prove that
$\left(M^{(q^{0},r^{0})}_{n}\right)_{n\in\mathbb{N}}:=\left(\sum_{k=1}^{n}H^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(H^{(q^{0},r^{0})}_{k}|\mathcal{F}_{n-1}\right)\right)_{n\in\mathbb{N}}$
is a martingale. We just have to prove that the sum is in $L^{1}.$ However, since every summand is in $L^{1},$ we immediately obtain the claim.
Now, we check the constraints of Theorem 3.2 in Hall and Heyde (2014) to prove that $M^{(q^{0},r^{0})}_{n}/\sigma_{n}\xRightarrow{n\to\infty}\mathcal{N}(0,1)$ for
$\sigma_{n}:=\sum_{k=1}^{n}\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}.$
Here, the notation does not represent that $\sigma_{n}$ depends on $q^{0},r^{0}.$ Finally, we infer the claim from this asymptotic distribution. All three constraints (3.18, 3.19 and 3.20), i.e.
$\max_{k}|\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}(\nabla D^{(q^{0},r^{0})}_{k})|/\sigma_{n}\xrightarrow[]{n\to\infty}_{p}0,$
$\sum_{k=1}^{n}\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}/\sigma_{n}^{2}\xrightarrow[]{n\to\infty}_{p}\eta^{2}=1,$
$\sup_{n}\mathbb{E}\left(\max_{k}\frac{\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}}{\sigma_{n}^{2}}\right)\leq c$
for $c\in\mathbb{N},$ follow directly by the uniformly boundedness of the transition and the emission probabilities of the hidden Markov chain. By using $\sigma^{(q^{0},r^{0})}_{n}/n$ convergences to the invertible matrix $J(q^{0},r^{0})$, we can directly infer the claim.
Finally, we are ready to prove the central limit theorem.
Proof of Theorem 2.
We proceed similar to Hoadley (1971). For simplicity, we write $\theta=(q,r)$. With probability $\epsilon_{M}\xrightarrow[]{M\to\infty}1$, it holds
$0=\nabla\ell((X_{1},...,X_{M}),\hat{\theta}^{M}).$
Hence, it holds
$\nabla\ell((X_{1},...,X_{M}),\theta)|_{\theta=\hat{\theta}^{M}}-\nabla\ell((X_{1},...,X_{M}),\theta^{0})$
$=(\hat{\theta}^{M}-\theta^{0})\cdot\int_{0}^{1}\nabla^{2}\ell((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)d\xi.$
We define $I_{M}(\theta^{0}):=\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)d\xi.$ Consequently,
$\sqrt{M}(\hat{\theta}^{M}-\theta^{0})I_{M}=\frac{1}{\sqrt{M}}\sum_{k=1}^{M}\nabla D_{k}((X_{1},...,X_{M}),\theta^{0}).$
We prove $I_{M}(\theta^{0})\xrightarrow[]{M\to\infty}_{p}J_{\theta^{0}}.$ It holds
$\lim_{M\to\infty}|I_{M}\left(\theta^{0}\right)-J_{\theta^{0}}|$
$=\lim_{M\to\infty}\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\left(\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)-\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0})\right)d\xi$
$=0,$
where we used dominated convergence. Then, with Proposition 5.12 and the assumption that the matrix $J_{\theta^{0}}$ is invertible to conclude the claim. ∎
6 Discussion
From a biological perspective, the Linkage Model is useful because it explains genetic data in a relatively simple framework while accounting for linkage. From a mathematical point of view, it is particularly interesting as its stationary distribution remains the same across all markers, even though the underlying Markov chain is inhomogeneous. This specific model has already been considered in the context of Markov chains (Saloff-Coste and Zúniga, 2007). In this work, we use this property to prove limit results for the MLE in the Admixture Model.
Specifically, we investigate the consistency and central limit theorems (CLTs) of the MLE in the Linkage Model. In doing so, we prove the uniqueness of the MLE in this model. This is an important result, as in the Admixture Model—even in the supervised setting—the MLE is sometimes non-unique (Pfaffelhuber and Rohde, 2022; Heinzel et al., 2025; Heinzel, 2025).
The theoretical results can be applied in several ways, for example in marker selection, similar to Pfaff et al. (2004), who proposed using CLTs for the Admixture Model. Marker selection remains a widely studied topic (Phillips et al., 2014; Kidd et al., 2014; Xavier et al., 2022, 2020; Pfaffelhuber et al., 2020; Resutik et al., 2023; Phillips et al., 2019; Kosoy et al., 2009), and the presented results can contribute to assessing the quality of a marker set. Furthermore, the CLT represents the first published approach to quantify the uncertainty of the MLE in the Linkage Model.
Arguably, the most important application of the CLT is that they provide a theoretical foundation for a statistical test to compare the Linkage Model with the Admixture Model. This test helps determine which model better fits a given dataset. To my knowledge, this is the first data-based model selection method for the Linkage Model, as recommended by Anderson and Burnham (2004).
Of course, both the Admixture Model and the Linkage Model are simplifications of biological reality. This study only addresses the question of which of the two models fits a dataset better. It is still an open problem whether either model describes the data appropriately. Future work could include goodness-of-fit tests to evaluate whether either model is adequate at all. It would also be interesting to compare the performance of different model selection methods, such as cross-validation (Anderson and Burnham, 2004), with the statistical test developed here.
In this study, we only considered the supervised setting, i.e., the allele frequencies are assumed to be known. However, in the unsupervised setting, determining the number of ancestral populations is a major challenge in population genetics (Evanno et al., 2005; Wang, 2019; Pritchard et al., 2000; Raj et al., 2014; Verity and Nichols, 2016; Alexander and Lange, 2011). So far, none of the existing methods perform well (Garcia-Erill and Albrechtsen, 2020). A promising approach for the Linkage Model—possibly also applicable to the Admixture Model—could be methods for choosing the order of a Hidden Markov Model (HMM), such as those proposed in van Handel (2008); MacKAY (2002).
Data availability
The python scripts (application of the statistical test, evaluation of the statistical test and calculation of the variance of the MLEs) are available on GitHub (https://github.com/CarolaHeinzel/LinkageModel).
Funding
CSH is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 499552394 – SFB Small Data.
Acknowledgment
CSH thanks Peter Pfaffelhuber for his continuous and strong support.
Declaration of Conflicts of Interest
The author declares that there are no conflicts of interest.

### v2

@@
-A Statistical Test for Comparing the Linkage and Admixture Model Based on Central Limit Theorems
+Highlights
 Abstract
@@
 The Linkage Model is a Hidden Markov Model (HMM) that extends the Admixture Model by incorporating linkage between neighboring loci.
-This study investigates the consistency and central limit behavior of maximum likelihood estimators (MLEs) for individual ancestry in the Linkage Model, complementing earlier results by (Pfaff et al., 2004; Pfaffelhuber and Rohde, 2022; Heinzel, 2025) for the Admixture Model.
-These theoretical results are used to prove theoretical properties of a statistical test that allows for model selection between the Admixture Model and the Linkage Model.
+We prove consistency and asymptotic normality of maximum likelihood estimators (MLEs) for the ancestry of individuals in the Linkage Model, complementing earlier results by (Pfaff et al., 2004; Pfaffelhuber and Rohde, 2022; Heinzel, 2025) for the Admixture Model.
+These results are used to prove that a statistical test that allows for model selection between the Admixture Model and the Linkage Model is an asymptotic level-$\alpha$-test.
 Finally, we demonstrate the practical relevance of our results by applying the test to real-world data from The 1000 Genomes Project Consortium (2015).
@@
 1 Introduction
-Both the Linkage Model (Falush et al., 2003) and the Admixture Model (e.g.
-Alexander et al.
-(2009)) are widely used to explain the genetic data of individuals.
-They are similar in that both assume that genetic data can be described by the ancestry proportions $q$ of an individual from $K$ ancestral populations and the allele frequencies in these populations.
-However, the Linkage Model additionally incorporates the distance between loci through a Hidden Markov Model (HMM).
-Furthermore, in the Linkage Model, there exists a parameter $r$, which can be interpreted as the number of generations since an admixture event (Falush et al., 2003).
-As a result, Falush et al.
-(2003) formulated the Linkage Model as a Hidden Markov Model (HMM).
+The Linkage Model (Falush et al., 2003) is widely used to explain the genetic data of individuals.
+It assumes that genetic data can be described by the ancestry proportions $q$ of an individual from $K$ ancestral populations and the allele frequencies in these populations.
+Furthermore, there exists a parameter $r$, which can be interpreted as the number of generations since an admixture event.
+The genetic distance between loci is also considered by using a Hidden Markov Model (HMM).
 The Admixture Model can be seen as a special case of the Linkage Model with $r=\infty$, i.e.
 the data is assumed to be independent across markers.
-A natural question that arises is which model fits a given dataset better.
-This leads to a nested model selection problem as described by Anderson and Burnham (2004).
-In our case, we consider the statistical hypothesis test:
+A natural question that arises is which model, the Admixture Model or the Linkage Model, fits a given dataset better.
+This leads to a nested model selection problem as described by e.g.
+Anderson and Burnham (2004).
+In our case, we consider the statistical hypothesis test
 $H_{0}:r=\infty\quad\text{vs.}\quad H_{1}:r\in[0,\infty).$ (1)
-This is a classical nested test problem, for which asymptotic theory exists (Wilks, 1938).
-However, the test is only valid as an asymptotic level-$\alpha$ test if the MLEs for both ancestry and $r$ are asymptotically normally distributed under the Linkage Model.
-Establishing asymptotic normality requires results from the theory of HMMs.
+This is a classical test problem, for which asymptotic theory exists (Wilks, 1938).
+However, the test is only valid as an asymptotic level-$\alpha$ test if the MLEs for both ancestry $q$ and $r$ are asymptotically normally distributed under the Linkage Model.
+Proving asymptotic normality requires results from the theory of HMMs.
 Consistency of MLEs in HMMs with finite state and observation spaces has been addressed in foundational work by Baum and Petrie (1966) and Petrie (1969).
@@
 (2005).
-However, the specific case of a time-inhomogeneous Markov chain that maintains the same stationary distribution across all time points—as occurs in the Linkage Model when $q$ is the initial distribution—has not yet been studied.
-Several statistical tests and heuristics have been proposed for model selection in genetics.
-For example, to decide whether a pair of loci is suitable for STRUCTURE (Pritchard et al., 2000), Kaeuffer et al.
+However, the specific case of a time-inhomogeneous Markov chain that maintains the same stationary distribution for the hidden chain across all time points—as occurs in the Linkage Model when $q$ is the initial distribution—has not yet been studied.
+The question whether the data should be used in STRUCTURE (Pritchard et al., 2000)(a software that estimates the ancestry and the allele frequencies for the Admixture Model or the Linkage Model) analysis has already been considered.
+For example, to decide whether a pair of loci is suitable for STRUCTURE, Kaeuffer et al.
 (2007) suggested using $r_{LD}$, a measure of linkage disequilibrium introduced by Hill and Robertson (1968).
 They ran STRUCTURE with the Linkage Model and used logistic regression to assess the impact of $r_{LD}$ on the detection of population structure.
-Other studies have examined the sensitivity of cluster identification to genetic and ecological constraints (Rosenberg et al., 2001; Evanno et al., 2005).
 To achieve our goal of establishing the theoretical properties of the test in (1), we proceed as follows: First, we precisely define both the Admixture Model and the Linkage Model.
-Next, we prove the asymptotic normality of the MLEs in the Linkage Model, both when the number of chromosomes and when the number of markers tends to infinity.
-Based on these results, we construct a statistical test for (1) and prove that it is an asymptotic level-$\alpha$ test.
-We also provide a method to quantify the uncertainty of the MLEs in the Linkage Model.
+Next, we prove the asymptotic normality of the MLEs in the Linkage Model, both when the number of observations tends to infinity.
+Based on these results, we quantify the uncertainty of the MLEs and we construct a statistical test for (1) and prove that it is an asymptotic level-$\alpha$ test.
 Finally, we evaluate the statistical test through simulations and apply it to real data from The 1000 Genomes Project Consortium (2015).
 2 Models
-We first define the Linkage Model for one individual and bi-allelic markers.
+We first define the Linkage Model for one individual and bi-allelic markers in Definition 2.1.
+The number of alleles on chromosome $c\in\{1,...,C\}$ at marker $m\in\{1,...,M_{c}\}$ is called $X_{c,m}\in\{0,1\}.$ We denote the genetic distance in centi Morgan (cM) between the loci m1m-1 and $m$ on chromosome $c$ by $d_{c,m}$.
+Recall that one centi Morgan corresponds to a $0.01$ crossing over between two genetic markers per Meiosis.
+Let $\mathbb{S}^{K}$ be the (K1)(K-1)-dimensional simplex and let $q:=(q_{1},\ldots,q_{K})\in\mathbb{S}^{K}$ be the ancestries of the individual from population 1,,K.1,...,K.
+The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, on chromosome $c$ is called $p_{c,k,m}.$ The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ on chromosome $c$.
+We write $(q^{0},r^{0})$ for the true values of $(q,r)$ and $\mathbb{E},\mathbb{P}$ for the expected value and the probability respectively, with respect to the true parameters $q^{0},r^{0}.$
 Definition 2.1 (Linkage Model for Haploid Individuals).
-The number of alleles at chromosome $c\in\{1,...,C\}$ at marker $m\in\{1,...,M_{c}\}$ is called $X_{c,m}\in\{0,1\}.$ We denote the genetic distance in centi Morgan (cM) between the loci m1m-1 and $m$ on chromosome $c$ by $d_{c,m}$.
-Let $\mathbb{S}^{K}$ be the (K1)(K-1)-dimensional simplex and let $q:=(q_{1},\ldots,q_{K})\in\mathbb{S}^{K}$ be the ancestries of the individual from population 1,,K.1,...,K.
-The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, chromosome $c$ is called $p_{c,k,m}.$ The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ at chromosome $c$.
-Based on this, we define the Markov process
+We define the Markov chain
 $\mathbb{P}_{q,r}(Z_{c,1}=k)$ $=q_{k},$
@@
 $\mathbb{P}_{q,r}(X_{c,m}=x|Z_{c,m})=\mathbb{P}_{q,r}\left(\textup{Ber}\left(q_{Z_{c,m}}p_{c,Z_{c,m},m}\right)=x\right).$
-We define the log-likelihood, for $M_{total}:=\sum_{c=1}^{M_{C}},$
+We define the log-likelihood, for $M_{total}:=\sum_{c=1}^{C}{M_{c}},$
 $\ell((x_{1,1},...,x_{C,M}),(q,r)):=\frac{\log\left(\mathbb{P}_{q,r}((X_{c,m}=x_{c,m})_{c=1,...,C,m=1,...,M_{c}}))\right)}{M_{total}}.$
-We write $(q^{0},r^{0})$ for the true parameters $(q,r)$ and $\mathbb{E},\mathbb{P}$ for the expected value and the probability respectively, with respect to the true parameters $q^{0},r^{0}.$
-We always assume that $d_{c,m}\geq 0$ for all $m\in\{1,\ldots,M\}$ and that the allele frequencies and $d_{c,m}$ are known.
+We additionally define the parameter space by $\Theta:=\{(q,r)\in[\kappa_{q},\kappa_{q}^{\prime}]^{K}\times[r_{lb},\infty]:\sum_{k=1}^{K-1}q_{k}=1\}\subset\mathbb{R}^{K+1}$.
+We always assume that the allele frequencies and $(d_{c,m})_{c=1,...,C,m=1,...,M_{c}}$ are known. Definition 2.1 is for haploid individuals, which can easily be extended to the diploid case.
 Remark 2.2 (Diploid Case).
 Let $Z^{j}:=(Z^{j}_{c,1},...,Z^{j}_{c,M}),j=1,2,$ be two independent Markov chains with transition matrix (2).
-Let $X_{c,m}^{dip}$ be the number of alleles at chromosome $c,$ marker $m$ and let $x\in\{0,1,2\}.$ The emission probabilities for the diploid case are defined by
+Let $X_{c,m}^{dip}$ be the number of alleles on chromosome $c,$ marker $m,$ for diploid individuals.
+The emission probabilities for the diploid case are defined by
 $\mathbb{P}_{q,r}\left(X^{dip}_{c,m}=x|Z^{1}_{c,m},Z^{2}_{c,m}\right)$
@@
 in The 1000 Genomes Project Consortium (2015).
-While phased data is important for the Linkage Model, the Admixture Model can deal with unphased data without any problems.
+While phased data is important for the Linkage Model, the Admixture Model can deal with unphased data without any problems. It is defined in Definition 2.3.
 Definition 2.3 (Admixture Model).
-The Admixture Model is a special case of the Linkage Model with $r=\infty.$
+The Admixture Model is a special case of the Linkage Model with $r=\infty.$ Especially, this means that it holds
+$X_{c,m}\sim\textup{Bin}(2,\langle q_{\cdot}^{0},p_{c,\cdot,m}\rangle).$
 Extending the models to a general number of individuals and markers with arbitrary number of alleles is straightforward.
-We do not only need to know about the Linkage Model, but we also need some general notation is mentioned in Notation 2.4.
-Notation 2.4 (General Notation).
-For a positive definite matrix $A$, we write $A\succ 0.$ Additionally, we write $X_{i:i+s}$ for $X_{i},...,X_{i+s}.$
+Finally, we define the statistical test that corresponds to test problem 2.
+Definition 2.4 (Statistical Test).
+We define the test statistic for the test (1) by
+$\Lambda:=-2\ln\left(\frac{\max\left\{\ell((X_{1,1},...,X_{C,M}),(q,\infty)):q\in\mathbb{S}^{K}\right\}}{\max\{\ell((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
+Furthermore, let $\chi^{2}_{1-\alpha}$ be the $1-\alpha$-quantile of the $\chi^{2}(1)$-distribution. We reject $H_{0}$, if $\Lambda>\chi^{2}_{1-\alpha}.$
 3 Theoretical Main Results
-In this chapter, we prove the main results, i.e. consistency and central limit results, if the number of markers, $M_{total},$ tends to infinity. Therefore, we need assumption 3.1.
+In this chapter, we prove the main results, i.e. consistency and central limit results, if the number of markers, $M_{total},$ tends to infinity. Therefore, we assume that Assumption 3.1 holds.
 Assumption 3.1.
-We assume that
-  • (A1)
-    the parameter space $\Theta:=\{[\kappa_{q},\kappa_{q}^{\prime}]^{K-1}\times[r_{lb},\infty]\}\subset\mathbb{R}^{K+1}$ is a compact set.
+We assume that $C=1$ and that we have haploid individuals. We additionally assume that
   • (A1)
@@
   • (A2)
-    $p_{k,m}\neq p_{\ell,m}$ for infinitely many markers m and for all $k,\ell\in\{1,..,K\},k\neq\ell.$
+    it holds $p_{k,m}\neq p_{\ell,m}$ for infinitely many markers m and for all $k,\ell\in\{1,..,K\},k\neq\ell.$
   • (A3)
-    Let $\lambda_{c,m}$ be the second largest of the transition matrix $(\mathbb{P}(Z_{c,m}=k|Z_{c,m-1}=k^{\prime}))_{k=1,...,K,k^{\prime}=1,...,K}$. It holds
-    $\prod_{c=1}^{C}\prod_{m=1}^{M_{c}}\lambda_{c,m}\xrightarrow[]{M_{total}\to\infty}0$
+    Let $\lambda_{c,m}$ be the second largest eigen value of the transition matrix $(\mathbb{P}(Z_{c,m}=k|Z_{c,m-1}=k^{\prime}))_{k=1,...,K,k^{\prime}=1,...,K}$. It holds
+    $\prod_{c=1}^{C}\prod_{m=1}^{M_{c}}\lambda_{c,m}\xrightarrow[]{M_{total}\to\infty}0.$
   • (A4)
     there exists a constant $\kappa_{d}>0$, such that $d_{c,m}\geq\kappa_{d}$ for all $m\in\{1,...,M_{c}\},c=1,...,C.$
+  • (A5)
+    It holds $d_{m}\leq d_{up}$ for all $m\in\{1,...,M\}.$
+By the definition of $\Theta$, we already assumed that $q\in[\kappa_{q},\kappa^{\prime}_{q}]$ and $r\geq r_{lb}.$
 Figure 1 gives an overview of the results in this paper.
-ConsistencyCLTAsymptotic Distribution of the Test Statistic
-Figure 1: Overview of results in this section and how they are connected to each other.
+Unique MLEAssumption 3.1ConsistencyAssumption 3.1CLTAssumption 3.1,invertible Fischer Information Asymptotic Distributionof the Test StatisticAssumption 3.1
+Figure 1: Overview of results in this section and their requirements.
 We also state the theory under the following conditions concerning the data.
@@
 Theorem 2 (Central Limit Theorem for the MLE).
-We define the Fischer information
-$J_{q^{0},r^{0}}:=\lim_{M_{total}\to\infty}\mathbb{E}\left(\frac{\partial^{2}}{\partial(q,r)^{2}}\ell\left((x_{1,1},...,x_{C,M}),(q,r))\big{|}_{(q,r)=\left(q^{0},r^{0}\right)}\right)\right)$
-and assume $J_{q^{0},r^{0}}\succ 0$. Let the true parameter be in the interior of the parameter space. Then, it holds
-$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow{M_{total}\to\infty}\mathcal{N}\left(0,J_{q^{0},r^{0}}^{-1}\right).$
+Let Assumption 3.1 holdand let the Fischer information
+$J(q^{0},r^{0}):=\lim_{M_{total}\to\infty}\mathbb{E}\left(\frac{\partial^{2}}{\partial(q,r)^{2}}\ell\left((x_{1,1},...,x_{C,M}),(q,r))\big|_{(q,r)=\left(q^{0},r^{0}\right)}\right)\right)$
+be positive definite. Let the true parameters be in the interior of the parameter space. Then, it holds
+$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow{M_{total}\to\infty}\mathcal{N}\left(0,J(q^{0},r^{0})^{-1}\right).$
 Remark 3.3.
@@
 in our case, if the allele frequencies and the distance between the markers are identical for every marker.
-This also means that the markers are all at different chromosomes or at the same chromosome.
-According to Wilks (1938), since the MLE in the Linkage Model is asymptotically normally distributed, Theorem 3 holds.
-Definition 3.4 (Statistical Test).
-We define the test statistic for the test (1) by
-$\Lambda:=-2\ln\left(\frac{\textup{argmax}\{(q,r)\mapsto\ell((X_{1,1},...,X_{C,M}),(q,\infty))\}}{\textup{argmax}\{(q,r)\mapsto\ell((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
-Furthermore, let $\chi^{2}_{1-\alpha}$ be the $1-\alpha$-quantile of the $\chi^{2}(1)$-distribution. We reject $H_{0}$, if $\Lambda>\chi^{2}_{1-\alpha}.$
+This also means that the markers are either all on different chromosomes or on the same chromosome.
+According to Wilks (1938), since the MLE in the Linkage Model is asymptotically normally distributed, Theorem 3 about the asymptotic distribution of the test statistic holds.
 Theorem 3 (Asymptotic distribution of the test statistic).
@@
 4 Application to Data
-In this chapter, we first evaluate the performance of the test from Definition 3.4 by using simulated data.
+In this section, we first evaluate the performance of the test from Definition 2.4 by using simulated data.
 Afterwards, we apply the test to data from The 1000 Genomes Project Consortium (2015) and compare the uncertainty of the MLEs in the Linkage Model to the ones in the Admixture Model.
 4.1 Simulated Data
-To evaluate the statistical test by using the type 1 and type 2 error, we simulate the data according to the Linkage Model and the Admixture Model with different values for $M,r$ and $(d_{c,m})_{c=1,...,C,m=1,...,M_{c}-1}.$ More precisely, we simulated 100 markers on one single chromosome with $d_{1,m}=d_{1,m+1}\forall m,d:=d_{1,m}\in\{0.1,0.5,1,2,5,10\}$ and $r\in\{1,10,100\}.$ We choose $0.05$ as the significance level.
-We calculated the type 1 and the type 2 error by repeating the experiment 100 times per possible combination of $r$ and $d$.
+To evaluate the statistical test in terms of Type 1 and Type 2 errors, I simulated data under both the Linkage Model and the Admixture Model for different values of $M$, $r$, and the marker distances $(d_{c,m})_{c=1,\dots,C,\ m=1,\dots,M_{c}-1}$.
+Specifically, I simulated 100 markers on a single chromosome with $d_{1,m}=d_{1,m+1}=:d$ for all $m\in\{1,...,M-1\}$, where $d\in\{0.1,0.5,1,2,5,10\}$ and $r\in\{1,10,100\}$.
+The significance level was set to $0.05$.
+I computed the Type 1 and Type 2 error rates by repeating the experiment 100 times for each possible combination of $r$ and $d$.
 The results are shown in Figure 2.
-Especially the type 1 error is small and lower than 0.05.
-Furthermore, for increasing $r$ and $d$ the type 1 and the type 2 error tend to increase.
-This is not surprising since a large $r$ corresponds to a high similarity between the two models.
-To summarize, the simulation results show that the test has both, a high power and a small type 1 error.
-Figure 2: Evaluation of the statistical test for simulated data.
+Notably, the Type 1 error remains small and consistently below the $0.05$ threshold.
+Moreover, both the Type 1 and Type 2 errors tend to increase with increasing values of $r$ and $d$.
+This behavior is expected, as larger values of $r$ and $d$ imply greater similarity between the two models, making them harder to distinguish.
+In summary, the simulation results indicate that the test exhibits high power while maintaining a low Type 1 error rate.
+Figure 2: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$.
 4.2 Real Data
-We also applied our theoretical results to the data from The 1000 Genomes Project Consortium (2015) with the AIM set by Kidd et al.
-(2014) that consists of $M_{total}=55$ markers on 20 different autosomes ($C=20$).
+I also applied the statistical test to the data from The 1000 Genomes Project Consortium (2015) with the ancestry informative marker (AIM) set by Kidd et al.
+(2014) consisting of $M_{total}=55$ markers on 20 different autosomes ($C=20$).
+For the analysis, individuals from five populations were included: Africa (AFR), Europe (EUR), South Asia (SAS), East Asia (EAS), and Admixed Americans (AMR), i.e., $K=5$.
 We always used all 2504 individual except the one that we are testing to calculate the allele frequencies as relative frequency.
@@
 (2014) in cM are shown in Figure 3.
-Figure 3: Genetic Distances of the markers in the AIM set by Kidd et al. (2014). We rounded them to integer values.
-The distance between the loci at a single chromosome was between below 1 cM and approximately 129 cM.
-We used individuals from Africa (AFR), Europe (EUR), South-East-Asian (SAS), East-Asian (EAS) and Admixed Americans (AMR), i.e.
-$K=5$.
+Figure 3: Genetic Distances of the markers in the AIM set by Kidd et al. (2014) rounded to integer values.
+The distance between loci on a single chromosome ranged from below 1 cM to approximately 129 cM.
 The results of the statistical test are shown in Figure 4.
-Altogether, for 312 out of 2504 ($\approx 12.5\%$ of the) individuals, the null hypothesis that the Admixture Model fits better to the data, cannot be rejected.
-Thereby, we see strong differences between the populations: For people from Africa and East Asian, the fraction of individuals for which the null hypothesis is rejected is much smaller than for the other populations.
-This emphasizes that it is not sufficient to name a general rule to decide for either the Linkage Model or the Admixture Model that only depends on the marker set.
-Instead, a method – like the statistical test (1) – is required, that decides for every individual or population separately.
-Figure 4: Results of the statistical test.
-We also compare the covariance for both, the MLE in the Linkage Model and the MLE in the Admixture Model to each other for individual HG00096 from Great Britain.
-Therefore, we again used the AIM Set by Kidd et al.
+In total, for 312 out of 2504 individuals (approximately 12.5%), the null hypothesis — that the Admixture Model provides a better fit to the data — cannot be rejected.
+Notably, there are substantial differences across populations: for individuals from Africa and East Asia, the proportion for which the null hypothesis is rejected is considerably lower compared to other groups.
+This highlights that it is not sufficient to define a general rule for choosing between the Linkage Model and the Admixture Model based solely on the marker set.
+Instead, a method — such as the statistical test defined in equation (1) — is needed, which allows model selection on an individual or population-specific basis.
+Figure 4: Results of the statistical test from Definition 2.4 for the data from The 1000 Genomes Project Consortium (2015).
+We further compare the covariance matrices of the MLEs obtained under the Linkage Model and the Admixture Model for individual HG00096 from Great Britain.
+For this analysis, I again used the AIM set introduced by Kidd et al.
 (2014).
-The results are presented in figures 6 and 6.
-There, we see that the variance of the estimators in the Linkage Model are higher than in the Admixture Model.
-This aligns with our expectations since the Linkage Model contains more parameters.
-Although both models clearly identify EUR as the primary ancestral population, the estimated values differ significantly.
+The results are shown in Figure 6 and Figure 6, corresponding to the Admixture and Linkage Models, respectively.
+As expected, the variance of the estimators is higher in the Linkage Model due to the larger number of parameters.
+While both models clearly identify EUR as the primary ancestral population, the estimated admixture proportions differ significantly between models (estimated ancestry of almost 1 from Europe in the Admixture Model and estimated ancestry of approximately 0.8 from Europe in the Linkage Model).
 Figure 5: Covariance Matrix for the MLE in the Admixture Model. We considered individual HG00096. The MLE for $q$ was $(1,0,0,0,0).$
@@
 We also provide the code for testing whether the Linkage Model or the Admixture Model fits better to the data of $N$ individuals on GitHub.
-Therefore, we could either assume that $r$ depends on the individual, i.e.
-we would have $NK$ parameters that we have to estimate or that we have only one global parameter $r,$ i.e.
-we estimate N(K1)+1N(K-1)+1 parameters.
-We decided for the latter case.
-However, the allele frequencies are supposed to be known in this case.
-Additionally, the diploid case is also implemented.
+In this framework, two modeling approaches are possible: either (i) the recombination rate $r$ is assumed to vary across individuals, requiring the estimation of $N\times K$ parameters, or (ii) a single global recombination rate $r$ is assumed, reducing the number of parameters to $N\times(K-1)+1$.
+We chose the latter approach (ii) to reduce model complexity.
 5 Proofs of the Main Results
-We first prove that the MLE is asymptotically unique.
+We first prove that the MLE $\left(\hat{Q}^{M},\hat{R}^{M}\right):=\left(\hat{Q}^{1,M},\hat{R}^{1,M}\right)$ is asymptotically unique.
 Based on this, we prove consistency and central limit results for the MLE in the Linkage Model.
@@
 5.1 Unique MLE
-In this section, we use that the limit
-$\ell((X_{1},...,X_{M}),(q,r))\xrightarrow[]{M\to\infty}\ell(q,r)$
-exists almost surely. We will prove this in section 5.2.
 Proving the uniqueness of a MLE for HMM has been already considered extensively for a finite state space, see e.g.
@@
 we aim to prove Theorem 4.
-Theorem 4.
+Therefore, we use that the limit
+$\ell((X_{1},...,X_{M}),(q,r))\xrightarrow[]{M\to\infty}\ell(q,r)$ (3)
+exists $\mathbb{P}$-almost surely. We will prove this in section 5.2.
+Theorem 4 (Unique Maximum Likelihood Estimator).
 Let Assumption 3.1 hold. Then, it holds
 $\ell(q^{0},r^{0})>\ell(q,r)\Leftrightarrow q,r\neq q^{0},r^{0}.$ (4)
-Therefore, we first prove that the parameters are identifiable based on a finite number of observations.
+To prove Theorem 4, we first prove that the parameters are identifiable based on a finite number of observations (Lemma 5.1).
 Based on this, we prove that if $\ell(q^{0},r^{0})\neq\ell(q,r)$, then the measures $\mathbb{P}_{q,r}|_{(X_{m})_{m=1,2,...}}$ and $\mathbb{P}_{q^{0},r^{0}}|_{(X_{m})_{m=1,2,...}}$ are mutually different.
-Based on this, we can then conclude the claim.
-We write $(q,r)\sim(q^{\prime},r^{\prime})$, if there exists an invariant distribution $\pi,\pi^{\prime}$ such that $\mathbb{P}_{(q,r)}^{X}=\mathbb{P}_{(q^{\prime},r^{\prime})}^{X}$ holds.
-Therefore, we only consider the case $K=2$ as the proof involves naming conditions that yield to the full rank of the emission matrix.
+Then, we can conclude the claim.
+We start by proving the identifiability of the parameter for $K=2.$
 Lemma 5.1 (Identifiability of the Parameters).
-Let Assumption 3.1 hold and let $K=2.$ If either the allele frequencies $p_{\cdot,i},...,p_{\cdot,i+s}$ are linearly independent and $s\geq 4$ or $s=3$ and $q\neq p_{1,2}/(p_{1,2}+p_{2,2})$ holds and in addition $p_{1,m}\neq p_{2,m}\forall m\in\{i,...,i+s\}$, the parameters $(q,r)$ are identifiable based on the random variables $X_{i},...,X_{i+s}.$
+Let $K=2$ and let $d_{1},d_{2}>0.$ If the allele frequencies of $p_{\cdot,i}$ and $p_{\cdot,i+2}$ are linearly independent and in addition $p_{1,m}\neq p_{2,m}\forall m\in\{i,i+1,i+2\}$, the parameters $(q,r)$ are identifiable based on the random variables $X_{i},...,X_{i+2}.$
 Proof.
@@
 $[M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}=\sum_{i=1}^{2}m^{1}_{i}(x_{1})m^{2}_{i}(x_{2})m^{3}_{i}(x_{3})$
-and $M_{1},M_{2},M_{3}]$ is a three-dimensional array for which the element $(x_{1},x_{2},x_{3})$ equals $M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}.$ It holds
-$\mathbb{P}(X_{1},X_{2},X_{3})$ $=\sum_{z_{1},z_{2},z_{3}}\pi_{z_{1}}A_{1}(z_{1},z_{2})A_{2}(z_{2},z_{3})B_{2}(X_{2},z_{2})B_{3}(X_{3},z_{3})B_{1}(X_{1},z_{1})$
-$=[\underbrace{\pi B_{1}A_{1}}_{=:M_{1}},B_{2},\underbrace{A_{2}B_{3}}_{=:M_{2}}].$
+and $M_{1},M_{2},M_{3}]$ is a three-dimensional array for which the element $(x_{1},x_{2},x_{3})$ equals $M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}.$ Without loss of generality, we assume $i=1$. It holds
+$\mathbb{P}(X_{1}$ $=x_{1},X_{2}=x_{2},X_{3}=x_{3})$
+$=\sum_{z_{1},z_{2},z_{3}}\pi_{z_{1}}A_{1}(z_{1},z_{2})A_{2}(z_{2},z_{3})B_{2}(x_{2},z_{2})B_{3}(x_{3},z_{3})B_{1}(x_{1},z_{1})$
+$=[\pi B_{1}A_{1},B_{2},\underbrace{A_{2}B_{3}}_{=:M_{2}}]_{x_{1},x_{2},x_{3}}.$
 The proof is divided into two steps:
   • (i)
-    prove that the entries of $M_{1},M_{2},B_{2}$ are identifiable from the marginal density of $X_{1},X_{2},X_{3}$. Therefore, we use theory by Allman et al. (2009); Kruskal (1977).
+    prove that the entries of $M_{1}:=B_{1}A_{1},M_{2},B_{2}$ are identifiable from the marginal density of $X_{1},X_{2},X_{3}$. Therefore, we use theory by Allman et al. (2009); Kruskal (1977).
   • (ii)
@@
 $\det(B_{2})=\bigg|\begin{pmatrix}q_{1}p_{1,2}&(1-q_{1})p_{2,2}\\ 1-q_{1}p_{1,2}&1-(1-q_{1})p_{2,2}\end{pmatrix}\bigg|=q_{1}p_{1,2}-(1-q_{1})p_{2,2}.$
-Hence, $B_{2}$ has the full rang if and only if $q_{1}\neq\frac{p_{1,2}}{p_{1,2}+p_{2,2}}.$ This is the reason, why we need the condition about the linear independence of the allele frequencies (since we can apply Kruskal’s theorem to two different time points within the interval).
-We have
+Hence, $B_{2}$ has the full rang if and only if $q_{1}\neq\frac{p_{1,2}}{p_{1,2}+p_{2,2}}.$ We have
 $M_{1}=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$
 with entries
-$a$ $=(1-q_{1}p_{1,1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+\left(1-(1-q_{1})p_{2,1}\right)(1-e^{-rd_{1}})(1-q_{1})$
+$a$ $=\mathbb{P}(X_{1}=0|Z_{2}=1)$
+$=\sum_{z_{1}=1}^{2}\mathbb{P}(X_{0}=0|Z_{1}=z_{1})\mathbb{P}(Z_{1}=z_{1}|Z_{2}=1)$
+$=(1-q_{1}p_{1,1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+\left(1-(1-q_{1})p_{2,1}\right)(1-e^{-rd_{1}})(1-q_{1})$
 $=(1-q_{1}p_{1,1})(q_{1}+e^{-rd_{1}}(1-q_{1}))+(1-p_{2,1}+q_{1}p_{2,1})(1-q_{1})(1-e^{-rd_{1}})$
-$b$ $=q_{1}p_{1,1}\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+p_{2,1}(1-q_{1})^{2}(1-e^{-rd_{1}}),$
-$c$ $=(1-q_{1}p_{1,1})(1-e^{-rd_{1}})(1-q_{1})+\left(1-(1-q_{1})p_{2,1}\right)\left((1-e^{-rd_{1}})(1-q_{1})+e^{-rd_{1}}\right),$
-$d$ $=q_{1}p_{1,1}(1-e^{-rd_{1}})(1-q_{1})+p_{2,1}(1-q_{1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right).$
+$b$ $=\mathbb{P}(X_{1}=1|Z_{2}=1)$
+$=q_{1}p_{1,1}\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+p_{2,1}(1-q_{1})^{2}(1-e^{-rd_{1}}),$
+$c$ $=\mathbb{P}(X_{1}=0|Z_{2}=2)$
+$=(1-q_{1}p_{1,1})(1-e^{-rd_{1}})(1-q_{1})+\left(1-(1-q_{1})p_{2,1}\right)\left((1-e^{-rd_{1}})(1-q_{1})+e^{-rd_{1}}\right),$
+$d$ $=\mathbb{P}(X_{1}=1|Z_{2}=2)$
+$=q_{1}p_{1,1}(1-e^{-rd_{1}})(1-q_{1})+p_{2,1}(1-q_{1})\left((1-e^{-rd_{1}})(1-q_{1})+e^{-rd_{1}}\right).$
 It holds
-$\det(M_{1})=e^{-rd_{1}}(1-e^{-rd_{1}})q_{1}(1-q_{1})(p_{1,1}+p_{2,1}-2p_{1,1}p_{2,1})=0$
-$\Leftrightarrow(r=\infty)\lor(r=0)\lor(q_{1}\in\{0,1\})\lor(p_{1,1}=p_{2,1}=0)\lor(p_{2,1}=p_{1,1}=1).$
-Here, we excluded every case except the case $r=\infty.$ This case leads to the Admixture Model which has already been discussed by Heinzel (2025). We additionally calculate
-$\mathbb{P}(X_{k-2}=x\mid Z_{k-1}=z)$ $=\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x,Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
-$\propto\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
-$\propto\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-1}=z^{\prime}\mid Z_{k-2}=z)$
+$\det(M_{1})=(2e^{-2rd_{1}}q_{1}^{2}-3e^{-2rd_{1}}q_{1}+e^{-2rd_{1}}-4e^{-rd_{1}}q_{1}^{2}+6e^{-rd_{1}}q_{1}-3e^{-rd_{1}}+2q_{1}^{2}-3q_{1}+1)$
+$\quad\quad\quad\quad\quad\quad\cdot(p_{1,1}q_{1}+p_{2,1}q_{1}-p_{2,1})=0$
+$\Leftrightarrow q_{1}=\frac{p_{2,1}}{p_{1,1}+p_{2,1}}$
+$\quad\quad\lor(2e^{-2rd_{1}}q_{1}^{2}-3e^{-2rd_{1}}q_{1}+e^{-2rd_{1}}-4e^{-rd_{1}}q_{1}^{2}+6e^{-rd_{1}}q_{1}-3e^{-rd_{1}}+2q_{1}^{2}-3q_{1}+1)=0.$
+It holds
+$2e^{-2rd_{1}}q_{1}^{2}-3e^{-2rd_{1}}q_{1}+e^{-2rd_{1}}-4e^{-rd_{1}}q_{1}^{2}+6e^{-rd_{1}}q_{1}-3e^{-rd_{1}}+2q_{1}^{2}-3q_{1}+1=0$
+$\Leftrightarrow q_{1}=\frac{e^{-rd_{1}}}{e^{-rd_{1}}-1}\lor q_{1}=\frac{e^{-rd_{1}}-3}{2(e^{-rd_{1}}-1)}.$
+Fortunately, it holds $\frac{e^{-rd_{1}}}{e^{-rd_{1}}-1}\leq 0$ and $\frac{e^{-rd_{1}}-3}{2(e^{-rd_{1}}-1)}\geq 3/2.$ Hence, we do not have to take care of these terms.
+We additionally calculate
+$\mathbb{P}(X_{k-2}$ $=x\mid Z_{k-1}=z)$
+$=\sum_{z^{\prime}=1}^{2}\mathbb{P}(X_{k-2}=x,Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
+$\propto\sum_{z^{\prime}=1}^{2}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
 $\propto\underbrace{\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=0)}_{=(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}}\cdot\mathbb{P}(Z_{k-1}=0\mid Z_{k-2}=z)$
 $\hskip 28.45274pt+\underbrace{\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=1)}_{=(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}}\cdot\underbrace{\mathbb{P}(Z_{k-1}=1\mid Z_{k-2}=z)}_{=1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})}$
-$\propto(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}\left(1_{z=0}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right)$
-$\hskip 28.45274pt+(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}\left(1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right).$
+$\propto(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}\left(1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right)$
+$\hskip 28.45274pt+(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}\left(1_{z=2}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right).$
 This leads to
@@
 with entries
-$a_{2}$ $=(1-q_{1}p_{1,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right)+(1-(1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}})$
-$b_{2}$ $=(q_{1}p_{1,3})\left(e^{-rd_{2}}+q_{1}(1-e^{-rd_{2}})\right)+((1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}}),$
-$c_{2}$ $=(1-q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})+(1-(1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right),$
-$d_{2}$ $=(q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})+((1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right).$
-Hence, it holds $rk(M_{2})=1\Leftrightarrow p_{2,3}=p_{1,3},q_{1}=0.5.$ Now, we come to step (ii), i.e.
+$a_{2}$ $=(1-q_{1}p_{1,3})\left(e^{-rd_{2}}+q_{1}(1-e^{-rd_{2}})\right)+(1-(1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}})$
+$b_{2}$ $=(q_{1}p_{1,3})\left(e^{-rd_{2}}+q_{1}(1-e^{-rd_{2}})\right)+(1-q_{1})p_{2,3}q_{1}(1-e^{-rd_{2}}),$
+$c_{2}$ $=(1-q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})$
+$\quad\quad+(1-(1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right),$
+$d_{2}$ $=q_{1}p_{1,3}(1-q_{1})(1-e^{-rd_{2}})+(1-q_{1})p_{2,3}\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right).$
+It holds
+$\det(M_{2})=e^{-rd_{2}}(p_{1,3}q_{1}+p_{2,3}q_{1}-p_{2,3})=0$
+$\Leftrightarrow r=\infty\lor q_{1}=\frac{p_{2,3}}{p_{1,3}+p_{2,3}}.$
+For $r=\infty,$ it holds $\mathbb{P}(X_{1}=x)=(q_{1}p_{1,1}+(1-q_{1})p_{2,1})^{x}(1-(q_{1}p_{1,1}+(1-q_{1})p_{2,1}))^{1-x}.$ Then, $q$ is identifiable if and only if $p_{2,1}\neq p_{1,1}.$
+Now, we come to step (ii), i.e.
 we have to ensure that the parameters are uniquely determined based on the entries.
 For the emission probabilities, this is a direct consequence of the assumption $p_{k,m}\neq p_{\ell,m}.$ Then, since $d_{m}>0,$ it is also immediate to infer the value $r$ from the other parameters.
-Remark 5.2 (Identifiability in the Admixture Model for $K=2$).
-Heinzel (2025) did not consider the identifiability for finitely many markers in the Admixture Model. For $K=2$, it holds
-$\mathbb{P}(X_{m}=x)$ $=\langle q,p_{\cdot,m}\rangle^{x}(1-\langle q,p_{\cdot,m}\rangle)^{1-x}$
-$=(q_{1}p_{1,m}+(1-q_{1})p_{2,m})^{x}(1-q_{1}p_{1,m}-(1-q_{1})p_{2,m})^{1-x}$
-$=\begin{cases}p_{2,m}+q_{1}(p_{1,m}-p_{2,m})=:c_{m},x=1\\ 1-p_{2,m}+q_{1}(-p_{1,m}+p_{2,m}),x=0\end{cases}.$
-This leads to
-$q_{1}=\frac{c_{m}-p_{2,m}}{p_{1,m}-p_{2,m}},$
-if $p_{1,m}=p_{2,m}.$ Hence, $q_{1}$ is uniquely determined if and only if $p_{1,m}=p_{2,m}.$
-Remark 5.3.
 In the latter, we will – without a proof – just assume that the parameters are identifiable from the marginal density of the observations.
@@
 Additionally, we then have to verify again that the matrices of the decomposition of the marginal density have the full rank under certain conditions.
-Now, we aim to infer (3) from the identifiability of the parameters.
-Therefore, we first prove a law of large numbers.
-Then, based on this, we conclude (3).
-This is done by using the following particularity of the hidden chain (Lemma 5.4) and that $(Z_{m})_{m=1,2,...}$ is uniformly ergodic (Lemma 5.6).
+Now, we aim to infer (4) from the identifiability. Therefore, we first prove a law of large numbers. Then, based on this, we conclude (4).
+To prove Theorem 4, we will use the following particularity of the hidden chain.
 Lemma 5.2 (Invariant Measure).
 For each $m\in\mathbb{N},$ the probability measure $q$ is the unique invariant measure for the transition matrix of $Z_{m}$.
+Additionally, we will use that the hidden chain is uniformly ergodic (Lemma 5.4), which is defined as follows.
 Definition 5.3 (Uniformly Ergodic).
@@
 for $\rho_{\theta}<1,c_{\theta}\in\mathbb{R}.$
-Lemma 5.6 (Uniformly Ergodicity for the Linkage Model).
+Lemma 5.4 (Uniformly Ergodicity in the Linkage Model).
 Under Assumption 3.1, $(Z_{m})_{m=1,2,...}$ is uniformly ergodic.
 Proof.
-This claim is derived from Theorem 3.3 in Saloff-Coste and Zúniga (2007). ∎
-Now, we are ready to name an LLN for the Linkage Model.
+According to Saloff-Coste and Zúniga (2007), Theorem 3.3, the convergence result holds from Lemma 5.2. ∎
 Lemma 5.5.
-Let Assumption 3.1 hold and let $f_{i}$ be uniformly bounded by $f_{\infty}$. Then, it holds
-$\mathbb{P}^{\nu}_{\theta}\left(\frac{1}{n}\left|\sum_{i=1,...,n}f_{i}(X_{i:i+s})-\pi(f_{i})\right|\geq\epsilon\right)\leq C\exp\left(-\frac{n}{8f_{\infty}}\right)$
-for a constant $C>0$ and for any probability measure $\nu.$
-Proof.
-Without loss of generality, we assume $\mathbb{E}_{\pi}(f_{i}(X_{i:i+s}))=0$.
-From Lemma 5.6, we know that the Markov chain $(Z_{m})_{m=1,2,...}$ is uniformly ergodic.
-The proof is similar to Theorem 14 in Douc (2005): We first write
-$\xi_{i,j}:=\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i+j},X_{0:i+j})+\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i+j-1},X_{0:i+j-1}).$
-Then, since
-$\sum_{j=0}^{s}\sum_{i=1}^{n}\xi_{i,j}$ $=\sum_{i=1}^{n}\sum_{j=0}^{s}\xi_{i,j}=\sum_{i=1}^{n}\xi_{i,s}-\xi_{i,0},$
-it holds
-$\sum_{i=1,...,n}f_{i}(X_{i:i+s})=\sum_{j=0}^{s}\sum_{i=1}^{n}\xi_{i,j}+\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i-1},X_{0:i-1}).$
-Now, we can apply Azuma-Hoeffding inequality, which leads to
-$\mathbb{P}^{\nu}\left(\bigg{|}\sum_{i=1}^{n}\xi_{i,j}\bigg{|}\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{8nf_{\infty}}\right).$
-Now, we note that $\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|X_{0:i-1},Z_{0:i-1})=F(Z_{i-1})$ for a certain function $F.$ We now prove that
-$\mathbb{P}\left(\bigg{|}\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|X_{0:i-1},Z_{0:i-1})\bigg{|}\geq t\right)2\leq\exp\left(-\frac{t^{2}}{8nf_{\infty}}\right).$
-This is Theorem 17 in Douc et al. (2011).
-To prove this claim, we apply Etemadi’s inequality and Bernstein’s inequality to the $\xi_{k}$ as defined in Lemma 18 in Douc et al.
-(2011).
-Neither of them requires that the $\xi_{k}$ are identically distributed.
-Additionally, they used their Proposition 19 to prove the claim.
-Note, however, that in our case we can highly simplify the proof as we have a discrete Markov chain with bounded transition probabilities.
-Especially, we do not even need their $\check{X}_{n}$ as in our case, their $d_{n}$ equals 1 for all $n\in\mathbb{N}.$
-Lemma 5.8.
-Assumption 3.1 implies that there exists a uniformly bounded function $h_{i}$ and an integer $s\in\mathbb{N}$ such that
-$(i)\quad\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{M\to\infty}1\quad\mathbb{P}_{q^{0},r^{0}}-a.s.,$
-$(ii)\quad\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{M\to\infty}0\quad\mathbb{P}_{q,r}-a.s.$
+Assumption 3.1 implies that there exists a function $h_{i}$ and an integer $s\in\mathbb{N}$ such that
+$(i)\quad\frac{1}{n}\sum_{i=1}^{n}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{n\to\infty}1\quad\mathbb{P}_{(q^{0},r^{0})}\,-a.s.$
+$(ii)\quad\frac{1}{n}\sum_{i=1}^{n}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{n\to\infty}0\quad\mathbb{P}_{(q,r)}\,-a.s.$
 for every $(q,r)\neq(q^{0},r^{0}).$
@@
 (2011).
-However, in our case $X$ is not a stationary sequence.
-We find an integer $s$ and bounded functions $h_{i},i=1,2,$..., such that $\mathbb{E}_{\pi}(h_{i}(X_{i},...,X_{i+s}))=1,\mathbb{E}_{\pi^{0}}(h_{i}(X_{i},...,X_{i+s}))=0$ according to the theorem from Hahn-Banach for $\pi\neq\pi^{0}$.
-This is a consequence of Lemma 5.1.
-More precisely, we choose
-$h_{i}(X_{i},...,X_{i+s})=1-\frac{f_{i}(X_{i},...,X_{i+s})-1}{\mathbb{E}_{\pi^{0}}(f_{i}(X_{i},...,X_{i+s}))-1}$
-with
-$f_{i}(X_{i},...,X_{i+s}):=\frac{\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})}{\mathbb{P}_{\pi}(X_{i},...,X_{i+s})}$
-for $\pi\neq\pi^{0}.$ The case $\mathbb{E}_{\pi^{0}}(f_{i})=1$ cannot occur, which follows from Cauchy-Schwarz.
-However, the densities are uniformly bounded according to Assumption 3.1, i.e.
-we can also choose $h_{i}$ uniformly bounded.
-Hence, we use Dolgopyat and Sarig (2023), Theorem 3.12, to prove (i).
-For (ii), We define
-$A_{M}:=\left\{X_{0:M}:\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})\geq\frac{1}{2}\right\}.$
-It holds
-$\mathbb{P}_{q,r}^{\nu}(X_{0:M}\in A_{M})$
-$\quad=\sum_{k=1}^{K}\mathbb{P}_{q,r}^{\nu}(X_{0:M}\in A_{M}|Z_{0}=k)\mathbb{P}_{q,r}^{\nu}(Z_{0}=k)$
-$\quad\leq\max_{k=1,...,K}\sup_{\textup{supp }\mu}\mathbb{P}^{\mu}_{q,r}\left(\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})\geq\frac{1}{2}\right)$
-$\quad\leq c_{1}\exp(-M/c_{2})$
-for some constant $c_{1},c_{2}>0.$ Here, we use Lemma 5.7 for the last inequality. ∎
-Finally, we use Lemma 5.8 to prove Theorem 4.
+The only difference is that in our case $X$ is not a stationary sequence.
+We find an integer $s$ and a bounded functions $h_{i},i=1,2,$..., such that $\mathbb{E}_{\pi}(h_{i}(X_{i},...,X_{i+s}))=1,\mathbb{E}_{\pi^{0}}(h_{i}(X_{i},...,X_{i+s}))=0$ according to the theorem from Hahn-Banach for $\pi\neq\pi^{0}$.
+More precisely, we choose $h_{i}=1-\frac{f_{i}-1}{\mathbb{E}_{\pi^{0}}(f_{i})-1}$ with $f_{i}:=\frac{\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})}{\mathbb{P}_{\pi}(X_{i},...,X_{i+s})}$ for $\theta\neq\theta^{0}$.
+The case $\mathbb{E}_{\pi^{0}}(f_{i})=1$ cannot occur according to Lemma 5.1.
+However, we have to ensure that there exists a constant $c_{q,r,q^{0},r^{0}}>0,$ such that $\mathbb{E}_{\pi^{0}}(f_{i})\geq 1+c_{q,r,q^{0},r^{0}}.$ It holds
+$\mathbb{E}_{\pi^{0}}(f_{i})$ $=1+\int\frac{(\mathbb{P}_{\pi}(X_{i},...,X_{i+s})-\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s}))^{2}}{\mathbb{P}_{\pi}(X_{i},...,X_{i+s})}dX_{i}....X_{i+s}$
+$\geq 1+\frac{1}{\kappa^{2(s+1)}}\int(\mathbb{P}_{\pi}(X_{i},...,X_{i+s})-\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s}))^{2}\,dX_{i}....X_{i+s}$
+$\geq 1+\frac{1}{\kappa^{2(s+1)}K^{s+1}}|\mathbb{P}_{\pi}(X_{i},....,X_{i+s})-\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})|_{TV}$
+for $\kappa=\kappa_{q}\lor\kappa_{p}.$ There exists a $\delta>0$, such that $|\mathbb{P}_{\pi}(X_{m}=x|Z_{m}=z)-\mathbb{P}_{\pi^{0}}(X_{m}=x|Z_{m}=z)|\geq\delta$ for all $z_{m},z,x,z_{m-1}$, the total variation distance between $\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})$ and $\mathbb{P}_{\pi}(X_{i},...,X_{i+s})$ is bounded below.
+The densities are uniformly bounded according to Assumption 3.1, i.e. $h_{i}$ is uniformly bounded.
+Hence, we proceed similar to the proof of Theorem 3.12 in Dolgopyat and Sarig (2023) to prove (i) and (ii).
+Dolgopyat and Sarig (2023) considered the convergence of $\sum_{i=1}^{n}f_{i}(Z_{i},Z_{i+1})$ under certain conditions.
+Here, we consider the convergence of $\sum_{i=1}^{n}f_{i}(g_{i}(Z_{i}),g_{i+1}(Z_{i+1}))$ under certain conditions.
+The only difference compared to the proof in Dolgopyat and Sarig (2023) is that we have to prove
+$\|\mathbb{E}(f_{i}(g_{i}(Z_{i}),g_{i+1}(Z_{i+1}))-\mathbb{E}(f_{i}(g_{i}(Z_{i}),g_{i+1}(Z_{i+1}))|Z_{j})\|_{2}\leq c\rho^{i-j}$
+for $i>j.$ However, this claim is a direct consequence of Lemma 5.4.
+We use Lemma 5.5 to prove the identifiably of the parameters. This is stated in the following theorem (Theorem 7.13 in van Handel (2008)).
 Proof of Theorem 4.
@@
 $\ell(q,r)=\ell(q^{0},r^{0})\Rightarrow\mathbb{P}_{q^{0},r^{0}}|_{(X_{m})_{m=1,2,...}}<<\mathbb{P}_{q,r}|_{(X_{m})_{m=1,2,...}}$.
-Since the number of observable states is finite, we immediately obtain this.
-Now, we prove that $\mathbb{P}_{q^{0},r^{0}}\neq\mathbb{P}_{q,r}$ implies that $\mathbb{P}_{q^{0},r^{0}},\mathbb{P}_{q,r}$ are mutually singular.
-This is Lemma 5.8.
-Hence, we can then infer the claim analogously to van Handel (2008) in the proof of Theorem 7.13.
-∎
+Therefore, we use Shiryaev (2016). Since the number of observable states is finite, the claim is a direct consequence of this.
+Now, we prove that $\mathbb{P}_{q^{0},r^{0}}\neq\mathbb{P}_{q,r}$ implies that $\mathbb{P}_{q^{0},r^{0}},\mathbb{P}_{q,r}$ are mutually singular. This is Lemma 5.5. ∎
 5.2 Consistency
@@
 Therefore, we follow the ideas by van Handel (2008), which are stated in the homogeneous case.
-First, we note that both the transition probabilities and the emission probabilities are Lipschitz continuous in the parameters given Assumption 3.1.
-This is a direct consequence of the mean value theorem applied to $r\mapsto|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|.$
-We already know that the true value is the unique maximum point of the likelihood according to Theorem 4. Therefore, we first prove
-$\sup_{(q,r)\in\Theta}|\ell((X_{1},...,X_{M}),(q,r))-\ell(q,r)|\xrightarrow[]{M\to\infty}0(*).$
-Based on this, we almost immediately get the claim of Theorem 1.
-Proving (*) is Lemma 5.11 whose proof is divided into three steps:
+First, we note that both the transition probabilities and the emission probabilities are Lipschitz continuous in the parameters.
+Lemma 5.6.
+Let $d_{m}\leq d_{up}$ for all $m\in\{1,...,M\}.$ There exists two constants $L_{e}\in\mathbb{R}$ and $L_{t}\in\mathbb{R}$ such that
+$|\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})-\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})|$ $\leq L_{e}|(q,r)-(q^{\prime},r^{\prime})|$
+$|\mathbb{P}_{q,r}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})-\mathbb{P}_{q^{\prime},r^{\prime}}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})|$ $\leq L_{t}|(q,r)-(q^{\prime},r^{\prime})|.$
+Proof.
+It holds
+$|\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})-\mathbb{P}_{q^{\prime},r^{\prime}}(X_{m}=x|Z_{m}=z_{m})|$ $=|q^{\prime}_{z_{m}}-q_{z_{m}}|p_{z_{m},m}.$
+Hence, we can choose $L_{e}=1.$ It also holds
+$|\mathbb{P}_{q,r}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})-\mathbb{P}_{q^{\prime},r^{\prime}}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})|$
+$=|(1-e^{-rd_{m}})q_{k}-(1-e^{-r^{\prime}d_{m}})q^{\prime}_{k}|$
+$=\big|(1-e^{-rd_{m}})(q_{k}-q^{\prime}_{k})+q^{\prime}_{k}(e^{-r^{\prime}d_{m}}-e^{-rd_{m}})\big|$
+$\leq|1-e^{-rd_{m}}|\;|q_{k}-q^{\prime}_{k}|+|q^{\prime}_{k}|\,|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|$
+$\leq|q_{k}-q^{\prime}_{k}|+|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|.$
+Moreover, by the mean value theorem there exists $\xi\in[r,r^{\prime}]$ such that
+$|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|=d_{m}\,|r-r^{\prime}|\,e^{-\xi d_{m}}\leq d_{m}\,|r-r^{\prime}|.$
+We already know that the true value is the unique maximum point of the likelihood according to Theorem 4. Hence, we first prove
+$\sup_{(q,r)\in\Theta}|\ell((X_{1},...,X_{M}),(q,r))-\ell(q,r)|\xrightarrow[]{M\to\infty}0$ (5)
+under Assumption 3.1. Based on this, we almost immediately get the claim of Theorem 1.
+Proving (5) is Lemma 5.9 whose proof is divided into three steps:
   1. 1)
-    Prove $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$ Since both, the transition probabilities and the emission probabilities are uniformly bounded below, we can infer (according to the majorant criterion)
-    $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$
+    Prove $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$
   2. 2)
-    Prove $\ell((X_{1},...,X_{M}),(q,r))$ convergences for $M\to\infty$ a.s. We call the limit $\ell(q,r).$ This is Lemma 5.9
+    Prove $\ell((X_{1},...,X_{M}),(q,r))$ converges for $M\to\infty$ a.s. We call the limit $\ell(q,r).$ This is Lemma 5.7
   3. 3)
-    Prove the claim ()(*). This is Lemma 5.10.
-Lemma 5.9 is Lemma 7.8 in van Handel (2008), which easily can be adapted to the inhomogeneous case.
+    Prove equation (5). This is Lemma 5.9.
+Lemma 5.7 and Lemma 5.8 are Lemma 7.8 and Lemma 7.9 in van Handel (2008), respectively, and can easily be adapted to the inhomogeneous case.
 Lemma 5.7.
@@
 $D_{k,\ell}^{q,r}:=\log\left(\int\mathbb{P}_{q,r}(X_{k}|Z_{k}=z)\mathbb{P}_{q,r}(Z_{k}=z|X_{k-1},...,X_{k-\ell})dz\right)$
-There exist constants $\epsilon\in(0,1),\alpha$ so that
+There exists a constant $\alpha$ so that
 $\sup_{k\in\mathbb{N}}|D_{k,\ell}^{q,r}-D_{k}^{q,r}|\leq\alpha(1-\epsilon^{2})^{\ell}$
 for all $\ell,k\in\mathbb{N}.$
-Lemma 5.10 is Lemma 7.9 in van Handel (2008).
 Lemma 5.8.
@@
 for all $(q,r),(q^{\prime},r^{\prime})\in\Theta.$
+Lemma 5.9 is stated for the homogeneous case as Proposition 7.5 in van Handel (2008).
 Lemma 5.9.
-Let Assumption 3.1 hold. Then, it holds
-$\sup_{q,r\in\Theta}|\ell_{M}(q,r)-\ell(q,r)|\xrightarrow[]{M\to\infty}0.$
+Let Assumption 3.1 hold. Then, equation (5) holds.
 Proof.
 For simplicity, we write $\theta=(q,r)$ and $\theta^{0}=\left(q^{0},r^{0}\right).$ We start chronologically:
+  1. 1)
+    Since both, the transition probabilities and the emission probabilities are uniformly bounded below, we infer (according to the majorant criterion)
+    $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$
   2. 2)
@@
     The details can be found in van Handel (2008).
-Let us mention that we first prove that $\ell(q,r)$ is also Lipschitz continuous.
+Let us mention that we first prove that $\ell(q,r)$ is also Lipschitz continuous by using Lemma 5.6.
 Then, the claim is a direct consequence of Lemma 5.8 and the compactness of $\Theta$.
@@
 Additionally, we know that the MLE is unique, which leads to the claim.
-The details are described by van Handel (2008).
+The details are described by van Handel (2008), Theorem 7.6.
 ∎
 5.3 Central Limit Results
-To prove Theorem 2, we first prove that the first derivative of $\ell((X_{1},...,X_{M}),(q^{0},r^{0}))$ is asymptotically normally distributed.
+To prove the asymptotic normality of the MLE (Theorem 2), we first prove that the first derivative of $\ell((X_{1},...,X_{M}),(q^{0},r^{0}))$ is asymptotically normally distributed.
+Based on this, we conclude the claim with similar techniques as in Hoadley (1971).
 Proposition 5.10.
 Let $J(q^{0},r^{0})\succ 0$ and let Assumption 3.1 hold. Then, it holds
-$\lim_{M\to\infty}\frac{\nabla\ell((X_{1},...,X_{M}),(q^{0},r^{0}))}{\sqrt{M}}\xRightarrow{M\to\infty}\mathcal{N}\left(0,J^{-1}\left(q^{0},r^{0}\right)\right).$
+$\lim_{M\to\infty}\frac{\nabla\ell((X_{1},...,X_{M}),(q^{0},r^{0}))}{\sqrt{M}}\xRightarrow{M\to\infty}\mathcal{N}\left(0,J^{-1}_{q^{0},r^{0}}\right).$
 Proof.
-Let $H^{(q^{0},r^{0})}_{k}:=\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)$. First, we prove that
-$\left(M^{(q^{0},r^{0})}_{n}\right)_{n\in\mathbb{N}}:=\left(\sum_{k=1}^{n}H^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(H^{(q^{0},r^{0})}_{k}|\mathcal{F}_{n-1}\right)\right)_{n\in\mathbb{N}}$
-is a martingale. We just have to prove that the sum is in $L^{1}.$ However, since every summand is in $L^{1},$ we immediately obtain the claim.
+The proof idea is the same as the one of Proposition 5 in (Douc and Matias, 2001). We first define
+$N_{k}^{(q^{0},r^{0})}:=\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)$.
+Let $\mathcal{F}_{n}$ be the sigma field that is generated by $\nabla D^{(q^{0},r^{0})}_{k},k=1,...,n$. We notice that
+$\left(M^{(q^{0},r^{0})}_{n}\right)_{n\in\mathbb{N}}:=\left(\sum_{k=1}^{n}N_{k}^{(q^{0},r^{0})}+\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{n}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{n}\right)|\mathcal{F}_{n-1}\right)\right)_{n\in\mathbb{N}}$
+is a martingale.
 Now, we check the constraints of Theorem 3.2 in Hall and Heyde (2014) to prove that $M^{(q^{0},r^{0})}_{n}/\sigma_{n}\xRightarrow{n\to\infty}\mathcal{N}(0,1)$ for
-$\sigma_{n}:=\sum_{k=1}^{n}\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}.$
-Here, the notation does not represent that $\sigma_{n}$ depends on $q^{0},r^{0}.$ Finally, we infer the claim from this asymptotic distribution. All three constraints (3.18, 3.19 and 3.20), i.e.
+$\sigma_{n}:=\sum_{k=1}^{n}\left(N_{k}^{(q^{0},r^{0})}\right)^{2}.$
+Here, the notation omits that $\sigma_{n}$ depends on $q^{0},r^{0}.$
+All three constraints (i.e. equations 3.18, 3.19 and 3.20 in Hall and Heyde (2014)), i.e.
 $\max_{k}|\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}(\nabla D^{(q^{0},r^{0})}_{k})|/\sigma_{n}\xrightarrow[]{n\to\infty}_{p}0,$
@@
 $\sup_{n}\mathbb{E}\left(\max_{k}\frac{\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}}{\sigma_{n}^{2}}\right)\leq c$
-for $c\in\mathbb{N},$ follow directly by the uniformly boundedness of the transition and the emission probabilities of the hidden Markov chain.
+for $c\in\mathbb{R},$ follow directly by the uniformly boundedness of the transition and the emission probabilities of the hidden Markov chain.
 By using $\sigma^{(q^{0},r^{0})}_{n}/n$ convergences to the invertible matrix $J(q^{0},r^{0})$, we can directly infer the claim.
-Finally, we are ready to prove the central limit theorem.
+Finally, we are ready to prove the asymptotic normality of $\hat{Q}^{M},\hat{R}^{M}$.
 Proof of Theorem 2.
@@
 $0=\nabla\ell((X_{1},...,X_{M}),\hat{\theta}^{M}).$
-Hence, it holds
+Hence, we get
 $\nabla\ell((X_{1},...,X_{M}),\theta)|_{\theta=\hat{\theta}^{M}}-\nabla\ell((X_{1},...,X_{M}),\theta^{0})$
@@
 We define $I_{M}(\theta^{0}):=\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)d\xi.$ Consequently,
-$\sqrt{M}(\hat{\theta}^{M}-\theta^{0})I_{M}=\frac{1}{\sqrt{M}}\sum_{k=1}^{M}\nabla D_{k}((X_{1},...,X_{M}),\theta^{0}).$
-We prove $I_{M}(\theta^{0})\xrightarrow[]{M\to\infty}_{p}J_{\theta^{0}}.$ It holds
-$\lim_{M\to\infty}|I_{M}\left(\theta^{0}\right)-J_{\theta^{0}}|$
-$=\lim_{M\to\infty}\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\left(\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)-\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0})\right)d\xi$
+$\sqrt{M}(\hat{\theta}^{M}-\theta^{0})I_{M}(\theta^{0})=\frac{1}{\sqrt{M}}\sum_{k=1}^{M}\nabla D_{k}((X_{1},...,X_{M}),\theta^{0}).$
+We prove $I_{M}(\theta^{0})\xrightarrow[]{M\to\infty}_{p}J\left(\theta^{0}\right).$ It holds
+$\lim_{M\to\infty}|I_{M}\left(\theta^{0}\right)-J\left(\theta^{0}\right)|$
+$=\lim_{M\to\infty}\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\left(\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)-\mathbb{E}\left(\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0})\right)\right)d\xi$
 $=0,$
-where we used dominated convergence. Then, with Proposition 5.12 and the assumption that the matrix $J_{\theta^{0}}$ is invertible to conclude the claim. ∎
+where we used dominated convergence. Then, with Proposition 5.10 and the assumption that the matrix $J\left(\theta^{0}\right)$ is invertible, we conclude the claim. ∎
 6 Discussion
 From a biological perspective, the Linkage Model is useful because it explains genetic data in a relatively simple framework while accounting for linkage.
-From a mathematical point of view, it is particularly interesting as its stationary distribution remains the same across all markers, even though the underlying Markov chain is inhomogeneous.
+From a mathematical point of view, it is particularly interesting as the stationary distribution of the hidden Markov chain remains the same across all markers, even though the underlying Markov chain is inhomogeneous.
 This specific model has already been considered in the context of Markov chains (Saloff-Coste and Zúniga, 2007).
-In this work, we use this property to prove limit results for the MLE in the Admixture Model.
-Specifically, we investigate the consistency and central limit theorems (CLTs) of the MLE in the Linkage Model.
+In this work, we use this property to prove limit results for the MLE in the Linkage Model.
+Specifically, we investigate the consistency and the asymptotic normality of the MLE in the Linkage Model.
 In doing so, we prove the uniqueness of the MLE in this model.
@@
 Furthermore, the CLT represents the first published approach to quantify the uncertainty of the MLE in the Linkage Model.
-Arguably, the most important application of the CLT is that they provide a theoretical foundation for a statistical test to compare the Linkage Model with the Admixture Model.
+Arguably, the most important application of the CLT is that it provides a theoretical foundation for a statistical test to compare the Linkage Model with the Admixture Model.
 This test helps determine which model better fits a given dataset.
-To my knowledge, this is the first data-based model selection method for the Linkage Model, as recommended by Anderson and Burnham (2004).
+To my knowledge, this is the first data-based model selection method for the Linkage Model.
 Of course, both the Admixture Model and the Linkage Model are simplifications of biological reality.
@@
 So far, none of the existing methods perform well (Garcia-Erill and Albrechtsen, 2020).
-A promising approach for the Linkage Model—possibly also applicable to the Admixture Model—could be methods for choosing the order of a Hidden Markov Model (HMM), such as those proposed in van Handel (2008); MacKAY (2002).
+A promising approach for the Linkage Model—possibly also applicable to the Admixture Model—could be methods for choosing the order of a HMM, such as those proposed in van Handel (2008); Mackay (2002).
 Data availability

### v3

@@
-Highlights
+arXiv:2509.12734v3[math.ST] 27 Feb 2026
 Abstract
@@
 Central limit theorems for MLEs in HMMs have also been developed (Bickel et al., 1998; Douc et al., 2004; Jensen and Petersen, 1999; Brouste and Kleptsyna, 2010).
-Overviews of the statistical theory of HMMs can be found in Ephraim and Merhav (2002); Cappé et al.
-(2005).
+Overviews of the statistical theory of HMMs can be found in Ephraim and Merhav (2002); cappé2005inference.
 However, the specific case of a time-inhomogeneous Markov chain that maintains the same stationary distribution for the hidden chain across all time points—as occurs in the Linkage Model when $q$ is the initial distribution—has not yet been studied.
@@
 Let $\mathbb{S}^{K}$ be the (K1)(K-1)-dimensional simplex and let $q:=(q_{1},\ldots,q_{K})\in\mathbb{S}^{K}$ be the ancestries of the individual from population 1,,K.1,...,K.
-The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, on chromosome $c$ is called $p_{c,k,m}.$ The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ on chromosome $c$.
+The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, on chromosome $c$ is called $p_{c,k,m}.$ We assume in this work that the allele frequencies are known, which is called supervised setting.
+The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ on chromosome $c$.
 We write $(q^{0},r^{0})$ for the true values of $(q,r)$ and $\mathbb{E},\mathbb{P}$ for the expected value and the probability respectively, with respect to the true parameters $q^{0},r^{0}.$
@@
 $\mathbb{P}_{q,r}(Z_{c,1}=k)$ $=q_{k},$
-$\mathbb{P}_{q,r}(Z_{c,m}=\tilde{k}|Z_{c,m-1}=k)$ $=\begin{cases}e^{-d_{c,m}r}+\left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ if }k=\tilde{k}\\ \left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ else},\end{cases}$ (2)
+$\mathbb{P}_{q,r}(Z_{c,m}=\tilde{k}|Z_{c,m-1}=k)$ $=\begin{cases}e^{-d_{c,m}r}+\left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ if }k=\tilde{k}\\ \left(1-e^{-d_{c,m}r}\right)q_{\tilde{k}},\textup{ else},\end{cases}$ (2)
 $\mathbb{P}_{q,r}(Z_{c+1,1}=k|Z_{c,M_{c}})$ $=q_{k}.$
 The emission probability is defined by
-$\mathbb{P}_{q,r}(X_{c,m}=x|Z_{c,m})=\mathbb{P}_{q,r}\left(\textup{Ber}\left(q_{Z_{c,m}}p_{c,Z_{c,m},m}\right)=x\right).$
+$\mathbb{P}_{q,r}(X_{c,m}=x|Z_{c,m}=k)={p_{c,k,m}^{x}(1-p_{c,k,m})^{1-x}}.$
 We define the log-likelihood, for $M_{total}:=\sum_{c=1}^{C}{M_{c}},$
 $\ell((x_{1,1},...,x_{C,M}),(q,r)):=\frac{\log\left(\mathbb{P}_{q,r}((X_{c,m}=x_{c,m})_{c=1,...,C,m=1,...,M_{c}}))\right)}{M_{total}}.$
-We additionally define the parameter space by $\Theta:=\{(q,r)\in[\kappa_{q},\kappa_{q}^{\prime}]^{K}\times[r_{lb},\infty]:\sum_{k=1}^{K-1}q_{k}=1\}\subset\mathbb{R}^{K+1}$.
+We additionally define the parameter space by $\Theta:=\{(q,r)\in[\kappa_{q},\kappa_{q}^{\prime}]^{K}\times[r_{lb},\infty]:\sum_{k=1}^{K-1}q_{k}{\leq}1\}\subset\mathbb{R}^{K+1}$.
+A visualization of the model is shown in Figure 1.
+[…]
+Figure 1: Visualization of the Linkage Model for two populations (blue and green) and three markers ($M=3$).
+(I): Reference Data Set that is used to calculate the allele frequencies.
+In this example we only consider bi-allelic markers.
+The letters A, C, G and T stand for the bases of the DNA, i.e.
+for Adenin, Cytosin, Guanin and Thymin.
+From this reference data set, we receive the allele frequencies, called $p_{k,m},k\in\{1,2\},m\in\{1,2,3\}.$ (II): Hidden Markov Model to describe the probabilities of the observed data $X_{1},X_{2},X_{3}.$ The hidden chain is represented as $Z_{1},Z_{2}$ and $Z_{3}$, the transition probabilities of state $Z_{i}$ to $Z_{i+1}$ are denoted as $A_{Z_{i},Z_{i+1}}$ and the emission probabilities $\mathbf{P}(X_{i}|Z_{i})$ are called $B_{Z_{i},X_{i}}.$ Ind.
+stands for Individual.
+(III): Estimators based on the reference data set and the Linkage Model.
 We always assume that the allele frequencies and $(d_{c,m})_{c=1,...,C,m=1,...,M_{c}}$ are known. Definition 2.1 is for haploid individuals, which can easily be extended to the diploid case.
@@
 The emission probabilities for the diploid case are defined by
-$\mathbb{P}_{q,r}\left(X^{dip}_{c,m}=x|Z^{1}_{c,m},Z^{2}_{c,m}\right)$
-$\hskip 28.45274pt=\begin{cases}q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}\cdot q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m},&x=2\\ q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}(1-q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m})+q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m}(1-q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}),&x=1\\ (1-q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m})(1-q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m}),&x=0.\end{cases}$
+${\mathbb{P}_{q,r}\left(X^{dip}_{c,m}=x|Z^{1}_{c,m}=z_{1},Z^{2}_{c,m}=z_{2}\right)}$
+$\hskip 28.45274pt{=}\begin{cases}{p_{c,z_{1},m}p_{c,z_{2},m},}&x=2\\ {p_{c,z_{1},m}(1-p_{c,z_{2},m})+p_{c,z_{2},m}(1-p_{c,z_{1},m}),}&x=1\\ {(1-p_{c,z_{1},m})(1-p_{c,z_{2},m})},&x=0.\end{cases}$
 For the diploid case, the information about the maternal and the paternal copies (Choi et al., 2018), i.e.
@@
 $X_{c,m}\sim\textup{Bin}(2,\langle q_{\cdot}^{0},p_{c,\cdot,m}\rangle).$
+A visualization of the Admixture Model is shown in Figure 2.
+[…]
+Figure 2: Visualization of the Admixture Model for two populations (blue and green) and three markers.
+(I): Reference Data Set that is used to calculate the allele frequencies.
+In this example we only consider bi-allelic markers.
+(II): Visualization of the model (III): Estimators based on the reference data set and the Admixture Model.
 Extending the models to a general number of individuals and markers with arbitrary number of alleles is straightforward.
@@
 We define the test statistic for the test (1) by
-$\Lambda:=-2\ln\left(\frac{\max\left\{\ell((X_{1,1},...,X_{C,M}),(q,\infty)):q\in\mathbb{S}^{K}\right\}}{\max\{\ell((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
+$\Lambda:=-2\ln\left(\frac{\max\left\{{\mathbb{P}_{q,r}}((X_{1,1},...,X_{C,M}),(q,\infty)):q\in\mathbb{S}^{K}\right\}}{\max\{{\mathbb{P}_{q,r}}((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
 Furthermore, let $\chi^{2}_{1-\alpha}$ be the $1-\alpha$-quantile of the $\chi^{2}(1)$-distribution. We reject $H_{0}$, if $\Lambda>\chi^{2}_{1-\alpha}.$
+The test in Definition 1 compares the relative support of the Linkage Model and the Admixture Model for the observed data.
+Importantly, this is a model-selection test: it implicitly assumes that one of the two models is adequate.
+It therefore does not address the possibility that both models provide a poor description of the data, in which case selecting the better model may still yield a misleading fit.
+This, however, is a different but widely considered research question as e.g.
+considered by (Tedeschi, 2006; Mimno et al., 2015; Carstens et al., 2022).
 3 Theoretical Main Results
@@
   • (A4)
-    there exists a constant $\kappa_{d}>0$, such that $d_{c,m}\geq\kappa_{d}$ for all $m\in\{1,...,M_{c}\},c=1,...,C.$
+    there exists a constant $\kappa_{d}>0$, such that $d_{m}\geq\kappa_{d}$ for all $m\in\{1,...,M\}.$
   • (A5)
-    It holds $d_{m}\leq d_{up}$ for all $m\in\{1,...,M\}.$
+    It holds $d_{m}\leq d_{up}$ for all but finitely many $m\in\{1,...,M\}.$
 By the definition of $\Theta$, we already assumed that $q\in[\kappa_{q},\kappa^{\prime}_{q}]$ and $r\geq r_{lb}.$
+Assumption 3.2 (Finitely many chromosomes).
+The assumption $C=1$ is imposed solely to simplify notation.
+The same theory extends to the general case $C\in\mathbb{N}$.
+Indeed, if the number of chromosomes is finite, then the limit $M\to\infty$ necessarily implies that there are infinitely many markers on at least one chromosome.
+Moreover, Assumption (A5) allows only finitely many exceptions to the condition $d_{m}\leq d_{\mathrm{ub}}$.
+In particular, there may be up to C1C-1 indices $m$ such that $d_{m}=\infty$, which is exactly the situation corresponding to $C$ distinct chromosomes.
 Figure 3 gives an overview of the results in this paper.
-Unique MLEAssumption 3.1ConsistencyAssumption 3.1CLTAssumption 3.1,invertible Fischer Information Asymptotic Distributionof the Test StatisticAssumption 3.1
+Unique MLEAssumption 3.1ConsistencyAssumption 3.1CLTAssumption 3.1,invertible FisherInformation Asymptotic Distributionof the Test StatisticAssumption 3.1
 Figure 3: Overview of results in this section and their requirements.
@@
 Theorem 2 (Central Limit Theorem for the MLE).
-Let Assumption 3.1 holdand let the Fischer information
+Let Assumption 3.1 hold and let the Fisher information
 $J(q^{0},r^{0}):=\lim_{M_{total}\to\infty}\mathbb{E}\left(\frac{\partial^{2}}{\partial(q,r)^{2}}\ell\left((x_{1,1},...,x_{C,M}),(q,r))\big|_{(q,r)=\left(q^{0},r^{0}\right)}\right)\right)$
 be positive definite. Let the true parameters be in the interior of the parameter space. Then, it holds
-$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow{M_{total}\to\infty}\mathcal{N}\left(0,J(q^{0},r^{0})^{-1}\right).$
+$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow[]{M_{total}\to\infty}\mathcal{N}\left(0,J(q^{0},r^{0})^{-1}\right).$
 Remark 3.4.
-Douc (2005) has proven the invertibility of the Fischer Information under certain constraints for stationary HMMs, i.e.
+Douc (2005) has proven the invertibility of the Fisher Information under certain constraints for stationary HMMs, i.e.
 in our case, if the allele frequencies and the distance between the markers are identical for every marker.
 This also means that the markers are either all on different chromosomes or on the same chromosome.
+Since there is no general theory about the invertibility of the Fisher Information yet, we as Dean et al.
+(2014), just assume that it is invertible.
+For the application of the test, this is not relevant.
 According to Wilks (1938), since the MLE in the Linkage Model is asymptotically normally distributed, Theorem 3 about the asymptotic distribution of the test statistic holds.
@@
 It holds
-$\Lambda\xRightarrow{M_{total}\to\infty}\chi^{2}(1)$
+$\Lambda\xRightarrow[]{M_{total}\to\infty}\chi^{2}(1)$
 under the null hypothesis of (1).
@@
 To evaluate the statistical test in terms of Type 1 and Type 2 errors, I simulated data under both the Linkage Model and the Admixture Model for different values of $M$, $r$, and the marker distances $(d_{c,m})_{c=1,\dots,C,\ m=1,\dots,M_{c}-1}$.
-Specifically, I simulated 100 markers on a single chromosome with $d_{1,m}=d_{1,m+1}=:d$ for all $m\in\{1,...,M-1\}$, where $d\in\{0.1,0.5,1,2,5,10\}$ and $r\in\{1,10,100\}$.
+Specifically, I simulated $M=50,100,200,500,1000,10000$ markers on a single chromosome with $d_{1,m}=d_{1,m+1}=:d$ for all $m\in\{1,...,M-1\}$, where $d\in\{0.1,1,2,10\}$ and $r\in\{0.1,1,10,\infty\}$.
+We evaluate the performance of the test for both, diploid and haploid individuals and for $K=2,5$.
 The significance level was set to $0.05$.
-I computed the Type 1 and Type 2 error rates by repeating the experiment 100 times for each possible combination of $r$ and $d$.
-The results are shown in Figure 2.
-Notably, the Type 1 error remains small and consistently below the $0.05$ threshold.
-Moreover, both the Type 1 and Type 2 errors tend to increase with increasing values of $r$ and $d$.
-This behavior is expected, as larger values of $r$ and $d$ imply greater similarity between the two models, making them harder to distinguish.
-In summary, the simulation results indicate that the test exhibits high power while maintaining a low Type 1 error rate.
-Figure 2: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$.
+I computed the Type 1 and Type 2 error rates by repeating the experiment 100 times for each possible combination of $r$ and $d$. The results are shown in Figure 4.
+Figure 4: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$ for $K=2$.
+The true ancestry was $q^{0}=(0.2,0.8).$ (A) $d_{m}=0.1$, (B) $d_{m}=1$, (C) $d_{m}=2$, (D) $d_{m}=10.$
+Figure 4 shows that the statistical test is a level-$\alpha$ test, even for $M=50$.
+Moreover, for small values of $r^{0}d$, the power of the test is high and, as expected, increases with $M$.
+However, for large values of $r^{0}d$, the power remains low even when $M=10000$.
+The reason is that $\exp(-r^{0}d)\approx 10^{-5}\approx 0$ when $r^{0}d=10$.
+In that case, we have
+$\mathbb{P}(X_{m}=x)$ $=\sum_{k=1}^{K}\mathbb{P}(X_{m}=x\mid Z_{m}=k)\mathbb{P}(Z_{m}=k)$
+$\approx\sum_{k=1}^{K}p_{k,m}q_{k}=\langle p_{\cdot,m},q_{\cdot}\rangle.$
+In particular, in this setting the test is unable to distinguish between the two models.
+Figure 5 shows the power of the statistical test for the same cases as in Figure 4, but for diploid individuals. We see that the power is slightly higher than in the haploid case.
+(B) Refer to caption
+Figure 5: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$ for $K=2$.
+We simulated diploid individuals.
+The true ancestry was $q^{0}=(0.2,0.8).$ (A) $d_{m}=0.1$, (B) $d_{m}=1$, (C) $d_{m}=2$, (D) $d_{m}=10.$
+Figure 6 shows the performance of the statistical test for $K=5$ in the haploid setting.
+Again, the results are comparable to those obtained for $K=2$ in the haploid case.
+This suggests that the value of $K$ does not have a substantial impact on the performance of the statistical test.
+Figure 6: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$ for $K=5$.
+The true ancestry was $q^{0}=(0.2,0.05,0.25,0.4,0.1).$ (A) $d_{m}=0.1$, (B) $d_{m}=1$, (C) $d_{m}=2$, D) $d_{m}=10.$
+The evaluation of the statistical test for $K=5$ and for diploid individuals are shown in Figure 15 in the Appendix.
+We also used msprime Baumdicker et al.
+(2022) to simulate data according to an island model with $K=5$ islands.
+We chose the 50, 100 or 1000 markers with the hightest allele frequency difference between the populations, i.e.
+with the hightest
+$\max\{p_{k}:k=1,...,5\}-\min\{p_{k}:k=1,...,5\}.$
+With the resulting allele frequencies and distances between the markers, we then simulate the data according to the Linkage or the Admixture Model, respectively.
+The results are shown in Figure 8 and Figure 8.
+Here, the power is much higher than in the previous simulations.
+The reason for this is that the values of $d$ are much smaller than in the simulation directly according to the model, which is shown in Figure 11 in the appendix 7.1.2.
+The migration rates were chosen to achieve $F_{ST}=0.05$.
+The appendix 7.1.2 contains a more detailed description of the simulation with msprime and for the simulations according to the model.
+Figure 7: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$ for $K=5$.
+Figure 8: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$ for $K=5$.
 4.2 Real Data
-I also applied the statistical test to the data from The 1000 Genomes Project Consortium (2015) with the ancestry informative marker (AIM) set by Kidd et al.
-(2014) consisting of $M_{total}=55$ markers on 20 different autosomes ($C=20$).
+I also applied the statistical test to the data from The 1000 Genomes Project Consortium (2015) with $M=350$ loci.
 For the analysis, individuals from five populations were included: Africa (AFR), Europe (EUR), South Asia (SAS), East Asia (EAS), and Admixed Americans (AMR), i.e., $K=5$.
-We always used all 2504 individual except the one that we are testing to calculate the allele frequencies as relative frequency.
-The values for $d$ for the AIM set by Kidd et al.
-(2014) in cM are shown in Figure 3.
-Figure 3: Genetic Distances of the markers in the AIM set by Kidd et al. (2014) rounded to integer values.
-The distance between loci on a single chromosome ranged from below 1 cM to approximately 129 cM.
+To select the loci, I first extracted every 2.000 SNP and then all bi-allelic loci, where both loci have at least $5\%$ frequency in every population.
+This then leads to 3340 markers.
+I uniformly chose $350$ markers out of these 3340 markers and apply the test to these 350 markers.
+The distances between the markers are shown in Figure 12 in the appendix.
+The code for this can be found on GitHub.
 The results of the statistical test are shown in Figure 9.
-In total, for 312 out of 2504 individuals (approximately 12.5%), the null hypothesis — that the Admixture Model provides a better fit to the data — cannot be rejected.
-Notably, there are substantial differences across populations: for individuals from Africa and East Asia, the proportion for which the null hypothesis is rejected is considerably lower compared to other groups.
-This highlights that it is not sufficient to define a general rule for choosing between the Linkage Model and the Admixture Model based solely on the marker set.
+In total, for 615 out of 2504 individuals (approximately 25%), the null hypothesis - that the Admixture Model provides a better fit to the data - cannot be rejected.
+Figure 9: Results of the statistical test from Definition 2.4 for the data from The 1000 Genomes Project Consortium (2015) with $M=350$.
+Interestingly, for EUR (38.77%), SAS (54%) or AMR (39.48%), the proportion of individuals for which the null hypothesis is rejected is much higher than for individuals from EAS (0.99%) or AFR (2.1%) .
+It might be that the linkage disequibrilium between the markers in these populations is higher in EAS and AFR or that the markers are less informative in these populations to detect linkage disequibrilium.
+However, these are only presumption as we cannot use the test to explain these differences between the populations.
+Note that we also applied the test with $M=3340$ markers.
+In this setting, the null hypothesis has to be rejected for all 2504 individuals.
+The results in Figure 9 indicates that it is not sufficient to define a general rule for choosing between the Linkage Model and the Admixture Model based solely on the marker set.
 Instead, a method — such as the statistical test defined in equation (1) — is needed, which allows model selection on an individual or population-specific basis.
-Figure 4: Results of the statistical test from Definition 2.4 for the data from The 1000 Genomes Project Consortium (2015).
-We further compare the covariance matrices of the MLEs obtained under the Linkage Model and the Admixture Model for individual HG00096 from Great Britain.
-For this analysis, I again used the AIM set introduced by Kidd et al.
-(2014).
-The results are shown in Figure 6 and Figure 6, corresponding to the Admixture and Linkage Models, respectively.
-As expected, the variance of the estimators is higher in the Linkage Model due to the larger number of parameters.
-While both models clearly identify EUR as the primary ancestral population, the estimated admixture proportions differ significantly between models (estimated ancestry of almost 1 from Europe in the Admixture Model and estimated ancestry of approximately 0.8 from Europe in the Linkage Model).
-Figure 5: Covariance Matrix for the MLE in the Admixture Model. We considered individual HG00096. The MLE for $q$ was $(1,0,0,0,0).$
-Figure 6: Covariance Matrix of the MLE in the Linkage Model. We considered individual HG00096. The MLE for $q$ was $(0.8,0.04,0.04,0.08,0.04)$ and for $r$ approximately $1.8.$
 Remark 4.1 (Test for a whole population).
-We also provide the code for testing whether the Linkage Model or the Admixture Model fits better to the data of $N$ individuals on GitHub.
+The same method is also applicable for testing whether the Linkage Model or the Admixture Model fits better to the data of $N$ individuals.
 In this framework, two modeling approaches are possible: either (i) the recombination rate $r$ is assumed to vary across individuals, requiring the estimation of $N\times K$ parameters, or (ii) a single global recombination rate $r$ is assumed, reducing the number of parameters to $N\times(K-1)+1$.
-We chose the latter approach (ii) to reduce model complexity.
+Remark 4.2 (Validity of Assumption 3.1 for the true data).
+We see in Figure 12 that Assumption (A4) and (A5) are met.
+The assumption (A2) is fulfilled for all markers and all possible combinations of the populations, while (A1) is also true due to the method how we chose the markers.
+However, e.g.
+for the $M=55$ marker in the AIM set by Kidd et al.
+(2014), this assumption is violated for 12 different $p_{k,m}.$ We need (A3) to justify the invertibility of the Fisher matrix, which is also true.
+The MLEs for $q$ was never equal to 0 or 1 and the MLE for $r$ was never equal to 0, i.e.
+also the assumption about the parameter space is met.
+Extended information about the real data set such as the precise allele frequencies and the MLEs can be found on GitHub, respectively.
+Remark 4.3 (Estimation of the Uncertainty).
+By comparing the two models to each other, the question which model has a lower uncertainty for the MLEs arises.
+This question is answered in Figure 14 and Figure 14 in the appendix by showing the covariance matrix of the MLEs.
+From this figures, we conclude that the uncertainty of the MLEs for $q^{0}$ is comparable small for the two models, while the uncertainty of $\hat{r}$ is relatively large.
 5 Proofs of the Main Results
@@
 We start with (i). Therefore, we have to prove $rk(M_{1})=rk(B_{2})=rk(M_{2})=2.$ It holds
-$\det(B_{2})=\bigg|\begin{pmatrix}q_{1}p_{1,2}&(1-q_{1})p_{2,2}\\ 1-q_{1}p_{1,2}&1-(1-q_{1})p_{2,2}\end{pmatrix}\bigg|=q_{1}p_{1,2}-(1-q_{1})p_{2,2}.$
-Hence, $B_{2}$ has the full rang if and only if $q_{1}\neq\frac{p_{1,2}}{p_{1,2}+p_{2,2}}.$ We have
+$\det(B_{2})=\bigg|\begin{pmatrix}p_{1,1}&1-p_{1,2}\\ p_{2,1}&1-p_{2,2}\end{pmatrix}\bigg|=p_{1,1}p_{2,2}-(1-p_{1,2})(1-p_{2,1}).$
+Hence, $B_{2}$ has the full rang if and only if $(p_{1,1},1-p_{1,2})$ and $(p_{2,1},1-p_{2,2})$ are linearly independent from each other. We have
 $M_{1}=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$
@@
 $a$ $=\mathbb{P}(X_{1}=0|Z_{2}=1)$
-$=\sum_{z_{1}=1}^{2}\mathbb{P}(X_{0}=0|Z_{1}=z_{1})\mathbb{P}(Z_{1}=z_{1}|Z_{2}=1)$
+$=\sum_{z_{1}=1}^{2}\mathbb{P}(X_{{1}}=0|Z_{1}=z_{1})\mathbb{P}(Z_{1}=z_{1}|Z_{2}=1)$
 $=(1-q_{1}p_{1,1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+\left(1-(1-q_{1})p_{2,1}\right)(1-e^{-rd_{1}})(1-q_{1})$
@@
 Lemma 5.6.
-Let $d_{m}\leq d_{up}$ for all $m\in\{1,...,M\}.$ There exists two constants $L_{e}\in\mathbb{R}$ and $L_{t}\in\mathbb{R}$ such that
-$|\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})-\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})|$ $\leq L_{e}|(q,r)-(q^{\prime},r^{\prime})|$
+Let $d_{m}\leq d_{up}$ for all $m\in\{1,...,M\}.$ There exists a constant $L_{t}:={\max\{1,d_{up}\}}\in\mathbb{R}$ such that
+$|\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})-\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})|$ $\leq|(q,r)-(q^{\prime},r^{\prime})|$
 $|\mathbb{P}_{q,r}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})-\mathbb{P}_{q^{\prime},r^{\prime}}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})|$ $\leq L_{t}|(q,r)-(q^{\prime},r^{\prime})|.$
@@
 It holds
-$|\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})-\mathbb{P}_{q^{\prime},r^{\prime}}(X_{m}=x|Z_{m}=z_{m})|$ $=|q^{\prime}_{z_{m}}-q_{z_{m}}|p_{z_{m},m}.$
-Hence, we can choose $L_{e}=1.$ It also holds
+$|\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})-\mathbb{P}_{q^{\prime},r^{\prime}}(X_{m}=x|Z_{m}=z_{m})|$ $={0}.$
+It also holds
 $|\mathbb{P}_{q,r}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})-\mathbb{P}_{q^{\prime},r^{\prime}}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})|$
@@
 Let $J(q^{0},r^{0})\succ 0$ and let Assumption 3.1 hold. Then, it holds
-$\lim_{M\to\infty}\frac{\nabla\ell((X_{1},...,X_{M}),(q^{0},r^{0}))}{\sqrt{M}}\xRightarrow{M\to\infty}\mathcal{N}\left(0,J^{-1}_{q^{0},r^{0}}\right).$
+$\lim_{M\to\infty}\frac{\nabla\ell((X_{1},...,X_{M}),(q^{0},r^{0}))}{\sqrt{M}}\xRightarrow[]{M\to\infty}\mathcal{N}\left(0,J^{-1}_{q^{0},r^{0}}\right).$
 Proof.
@@
 is a martingale.
-Now, we check the constraints of Theorem 3.2 in Hall and Heyde (2014) to prove that $M^{(q^{0},r^{0})}_{n}/\sigma_{n}\xRightarrow{n\to\infty}\mathcal{N}(0,1)$ for
+Now, we check the constraints of Theorem 3.2 in Hall and Heyde (2014) to prove that $M^{(q^{0},r^{0})}_{n}/\sigma_{n}\xRightarrow[]{n\to\infty}\mathcal{N}(0,1)$ for
 $\sigma_{n}:=\sum_{k=1}^{n}\left(N_{k}^{(q^{0},r^{0})}\right)^{2}.$
@@
 By using $\sigma^{(q^{0},r^{0})}_{n}/n$ convergences to the invertible matrix $J(q^{0},r^{0})$, we can directly infer the claim.
+To prove Theorem 2, we need to change limes and integral at some point, for which we need to prove the boundedness of the second derivation of $D_{m}.$
+Lemma 5.11 (Uniform boundedness of second derivatives of $D_{m}$).
+Let Assumption 3.1 hold. There exists a constant $C<\infty$ such that
+$\sup_{m\geq 1}\ \sup_{(q,r)\in\Theta}\left|\frac{\partial^{2}D_{m}(q,r)}{\partial\theta_{a}\,\partial\theta_{b}}\right|\leq C$
+for every combination $\theta_{a},\theta_{b}\in(q_{1},...,q_{k-1},r).$
+Proof.
+We remind ourselves on
+$D_{m}:=\log\left(\int\mathbb{P}_{q,r}(X_{m}|Z_{m}=z)\mathbb{P}_{q,r}(Z_{m}=dz|X_{1},...,X_{m-1})\right).$
+Hence, to prove the claim, we need to ensure that
+  • (i)
+    it holds $\inf_{m}\int\mathbb{P}_{q,r}(X_{m}|Z_{m}=z)\mathbb{P}_{q,r}(Z_{m}=dz|X_{1},...,X_{m-1})>c>0$ for some constant $c.$
+  • (ii)
+    the enumerator of $\frac{\partial^{2}D_{m}(q,r)}{\partial\theta_{a}\,\partial\theta_{b}}$ is also bounded.
+We start by proving (i): Let
+$b_{m,k}:=\mathbb{P}(X_{m}=x\mid Z_{m}=k),\qquad\lambda_{m}:=e^{-d_{m}r},$
+$\pi_{m,k}:=\mathbb{P}(Z_{m}=k\mid X_{1:m-1}),\qquad f_{m}:=\sum_{k=1}^{K}b_{m,k}\pi_{m,k},\qquad D_{m}(q,r):=\log f_{m}.$
+We start by proving $0<f_{m}<1$ for all $m.$ Set
+$b_{-}:=\min\{\kappa_{p},\ 1-\kappa_{p}^{\prime}\}>0,\qquad b_{+}:=\max\{\kappa_{p}^{\prime},\ 1-\kappa_{p}\}<1.$
+Since $x\in\{0,1\}$ and $\kappa_{p}\leq p_{k,m}\leq\kappa_{p}^{\prime}$, we have $b_{m,k}\in[b_{-},b_{+}]$ for all $m,k$.
+Because $(\pi_{m,k})_{k=1}^{K}$ is a probability vector, $f_{m}=\sum_{k}b_{m,k}\pi_{m,k}$ is a convex combination of numbers in $b_{-},b_{+}]$, hence
+$b_{-}\leq f_{m}\leq b_{+}\qquad\text{for all }m\geq 1.$ (6)
+We now continue with the proof of (ii). We therefore first need the uniform positivity of $\pi_{m,k}$. Since $\kappa_{d}\leq d_{m}$ and $r\geq r_{\mathrm{lb}}$, we have
+$\lambda_{m}=e^{-d_{m}r}\leq e^{-\kappa_{d}r_{\mathrm{lb}}}=:\lambda_{\max}<1,\qquad\Rightarrow\qquad 1-\lambda_{m}\geq 1-\lambda_{\max}=:c_{\lambda}>0.$
+For $m\geq 2$,
+$\pi_{m,k}=\lambda_{m}\alpha_{m-1,k}+(1-\lambda_{m})q_{k}\geq(1-\lambda_{m})q_{k}\geq c_{\lambda}\,\kappa_{q}=:\pi_{-}>0.$
+For $m=1$ $\pi_{1,k}=q_{k}\geq\kappa_{q}$, so the same type of lower bound holds. Thus, uniformly in $m$ and the parameters,
+$\pi_{m,k}\geq\pi_{-}>0\qquad\forall m,k.$ (7)
+The filtering update is
+$\alpha_{m,k}=\mathbb{P}(Z_{m}=k\mid X_{1:m})=\frac{b_{m,k}\pi_{m,k}}{f_{m}}.$
+Using (6), (7), and $b_{m,k}\geq b_{-}$ yields
+$\alpha_{m,k}\geq\frac{b_{-}\,\pi_{-}}{b_{+}}=:a_{-}>0\qquad\forall m,k.$
+Let $u:=(1/K,\dots,1/K)$. Then $\alpha_{m,k}\geq a_{-}$ implies
+$\alpha_{m}\ \geq\ \beta\,u\quad\text{componentwise, with}\quad\beta:=Ka_{-}\in(0,1].$
+Consequently, the filtering map $F_{m}$ (mapping $\alpha_{m-1}\mapsto\alpha_{m}$) satisfies a uniform total-variation contraction: there exists $\rho:=1-\beta\in[0,1)$ such that for all probability vectors $\alpha,\tilde{\alpha}$,
+$\|F_{m}(\alpha;q,r)-F_{m}(\tilde{\alpha};q,r)\|_{1}\leq\rho\,\|\alpha-\tilde{\alpha}\|_{1},$ (8)
+uniformly in $m$ and the parameters. In particular,
+$\sup_{m\geq 1}\sup_{(q,r)}\sup_{\alpha}\left\|\frac{\partial F_{m}}{\partial\alpha}(\alpha;q,r)\right\|_{1\to 1}\leq\rho<1.$ (9)
+We now prove uniform bounds for the second derivation of $\alpha_{m}$. For any free parameter $\theta\in\{q_{1},\dots,q_{K-1},r\}$ define
+$J_{m}^{(\theta)}:=\frac{\partial\alpha_{m}}{\partial\theta}\in\mathbb{R}^{K}.$
+By the chain rule for $\alpha_{m}=F_{m}(\alpha_{m-1};q,r)$,
+$J_{m}^{(\theta)}=\frac{\partial F_{m}}{\partial\alpha}(\alpha_{m-1};q,r)\,J_{m-1}^{(\theta)}+\frac{\partial F_{m}}{\partial\theta}(\alpha_{m-1};q,r).$
+All entries of $\partial F_{m}/\partial\theta$ are rational expressions in $b_{m,k},\pi_{m,k},f_{m}$ and in $\lambda_{m},\lambda_{m}^{\prime}$ (for $\theta=r$), with denominators bounded away from 0 by (6)–(7) and numerators uniformly bounded because $b_{m,k}\in[b_{-},b_{+}]$, $q_{k}\in[\kappa_{q},\kappa_{q}^{\prime}]$, $d_{m}\in[\kappa_{d},d_{\mathrm{up}}]$, and
+$\lambda_{m}\leq\lambda_{\max},\qquad|\lambda_{m}^{\prime}|=d_{m}\lambda_{m}\leq d_{\mathrm{up}}\lambda_{\max}.$
+Hence there exists $B_{1}<\infty$ such that
+$\sup_{m\geq 1}\sup_{(q,r)}\sup_{\alpha}\left\|\frac{\partial F_{m}}{\partial\theta}(\alpha;q,r)\right\|_{1}\leq B_{1}.$
+Together with (9), this yields
+$\|J_{m}^{(\theta)}\|_{1}\leq\rho\,\|J_{m-1}^{(\theta)}\|_{1}+B_{1},$
+and therefore $\sup_{m\geq 1}\|J_{m}^{(\theta)}\|_{1}<\infty$ by geometric summation.
+Similarly, for free parameters $\theta_{a},\theta_{b}$ define
+$H_{m}^{(\theta_{a},\theta_{b})}:=\frac{\partial^{2}\alpha_{m}}{\partial\theta_{a}\,\partial\theta_{b}}\in\mathbb{R}^{K}.$
+A second-order chain rule gives
+$H_{m}^{(\theta_{a},\theta_{b})}=\frac{\partial F_{m}}{\partial\alpha}\,H_{m-1}^{(\theta_{a},\theta_{b})}+\frac{\partial^{2}F_{m}}{\partial\alpha^{2}}\!\big[J_{m-1}^{(\theta_{a})},J_{m-1}^{(\theta_{b})}\big]+\frac{\partial^{2}F_{m}}{\partial\alpha\,\partial\theta_{a}}\,J_{m-1}^{(\theta_{b})}+\frac{\partial^{2}F_{m}}{\partial\alpha\,\partial\theta_{b}}\,J_{m-1}^{(\theta_{a})}+\frac{\partial^{2}F_{m}}{\partial\theta_{a}\,\partial\theta_{b}}.$
+All second derivatives of $F_{m}$ are uniformly bounded for the same reason as above, now also using
+$|\lambda_{m}^{\prime\prime}|=d_{m}^{2}\lambda_{m}\leq d_{\mathrm{up}}^{2}\lambda_{\max},$
+and the already established uniform bounds on $J_{m-1}$. Hence there exists $B_{2}<\infty$ such that
+$\|H_{m}^{(\theta_{a},\theta_{b})}\|_{1}\leq\rho\,\|H_{m-1}^{(\theta_{a},\theta_{b})}\|_{1}+B_{2},$
+which implies $\sup_{m\geq 1}\|H_{m}^{(\theta_{a},\theta_{b})}\|_{1}<\infty$.
+Based on the previous steps, we can conclude the uniform boundedness of $\nabla^{2}D_{m}$.
+Since $\pi_{m}=\lambda_{m}\alpha_{m-1}+(1-\lambda_{m})q$, the bounds on $\alpha_{m-1}$, $J_{m-1}$, $H_{m-1}$ together with uniform bounds on $\lambda_{m},\lambda_{m}^{\prime},\lambda_{m}^{\prime\prime}$ imply that $\pi_{m,k}$ and its first and second derivatives with respect to any free parameters are uniformly bounded in $m$.
+Therefore $f_{m}=\sum_{k}b_{m,k}\pi_{m,k}$ and its first/second derivatives are uniformly bounded as well.
+Finally,
+$\frac{\partial^{2}D_{m}}{\partial\theta_{a}\,\partial\theta_{b}}=\frac{f_{m}f_{m,ab}-f_{m,a}f_{m,b}}{f_{m}^{2}},\qquad f_{m,a}=\sum_{k}b_{m,k}\pi_{m,k,a},\quad f_{m,ab}=\sum_{k}b_{m,k}\pi_{m,k,ab},$
+and $f_{m}^{2}\geq b_{-}^{2}>0$ by (6). Hence the right-hand side is uniformly bounded in $m$ and the parameters, which proves the claim. ∎
 Finally, we are ready to prove the asymptotic normality of $\hat{Q}^{M},\hat{R}^{M}$.
@@
 $=0,$
-where we used dominated convergence. Then, with Proposition 5.10 and the assumption that the matrix $J\left(\theta^{0}\right)$ is invertible, we conclude the claim. ∎
+where we used dominated convergence.
+Therefore, we have to ensure that $\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0})$ is bounded a.s., which is Lemma 5.11.
+Then, with Proposition 5.10 and the assumption that the matrix $J\left(\theta^{0}\right)$ is invertible, we conclude the claim.
+∎
+Remark 5.12 (Extension to diploid data).
+To extend our proofs to general ploidy, especially to diploid data, we would have to deal with the uniqueness of the MLE in this case and we would also have to extend the assumption about the invertibility of the Fisher information to the diploid case.
+In particular, the proof ideas remain the same.
 6 Discussion
@@
 A promising approach for the Linkage Model—possibly also applicable to the Admixture Model—could be methods for choosing the order of a HMM, such as those proposed in van Handel (2008); Mackay (2002).
+Our theoretical results only hold for $M\to\infty$.
+However, there are many attempts to increase the power of the statistical test for smaller values of $M.$ This included bootstrap with Edgeworth expansion Hall (1992) or the Bartlett correction (Lawley, 1956; Hayakawa, 1977; Cordeiro, 1987).
+However, the theoretical statements only work for i.i.d.
+random variables here.
+Extending this theory to the Linkage Model and applying it to our test problem (1) might increase the power of the statistical test.
 Data availability
@@
 Acknowledgment
-CSH thanks Peter Pfaffelhuber for his continuous and strong support.
+CSH thanks two anonymous referees, who significantly improved the representation of the result.
 Declaration of Conflicts of Interest